### v2

# Behind the Myth of Exploration in Policy Gradients

Abstract
Policy-gradient algorithms are effective reinforcement learning methods for solving control problems. To compute near-optimal policies, it is essential in practice to include exploration terms in the learning objective. Although the effectiveness of these terms is usually justified by an intrinsic need to explore environments, we propose a novel analysis with the lens of numerical optimization. Two criteria are introduced on the learning objective and two others on its stochastic gradient estimates, and are afterwards used to discuss the quality of the policy after optimization. The analysis sheds the light on two separate effects of exploration techniques. First, they make it possible to smooth the learning objective and to eliminate local optima while preserving the global maximum. Second, they modify the gradient estimates, increasing the probability that the stochastic parameter updates eventually provide an optimal policy. These effects are illustrated empirically on exploration strategies based on entropy bonuses, highlighting their limitations and opening avenues for future works in the design and analysis of such strategies.
Reinforcement Learning, Policy Gradients, Exploration
1 Introduction
Many practical problems require making sequential decisions in environments. Reinforcement learning (RL) is a framework for solving such decision-making problems that has been successful on complex tasks, including playing games (Mnih et al., 2015; Silver et al., 2017), operating power systems (Aittahar et al., 2024), controlling robots (Kalashnikov et al., 2018), or interacting with electricity markets (Boukas et al., 2021).
Reinforcement learning algorithms interact with an environment to gather information about this environment, which in turn enables to compute and follow an optimal policy. This creates a trade-off between exploration and exploitation. In short, in order to eventually compute a good policy, it is necessary to obtain additional information about the environment by taking actions that are likely not optimal. In algorithms where the trade-off is explicit, exploration is well-understood and has been the subject of many works (Dann et al., 2017; Azar et al., 2017; Neu & Pike-Burke, 2020). In policy-gradient algorithms, one can most often not distinguish exploration from exploitation. Nevertheless, a main theoretical requirement to converge towards globally (or even locally) optimal solutions is that policies remain sufficiently stochastic during the learning procedure (Bhandari & Russo, 2019; Bhatt et al., 2019; Agarwal et al., 2020; Zhang et al., 2021a; Bedi et al., 2022). Interestingly, neither softmax nor Gaussian policies guarantee enough stochasticity for ensuring (fast) convergence (Mei et al., 2020, 2021; Bedi et al., 2022). This requirement of stochasticity in policy-gradient algorithms is often abusively called exploration and understood as the need to infinitely sample all states and actions.
Practitioners have tried to meet the theoretical requirement of sufficient randomness of policies in policy gradient via reward-shaping strategies, whereby a learning objective that promotes or hinders behaviors by providing reward bonuses for some states and actions is optimized as a surrogate to the return of the policy. These bonuses typically promote actions that reduce the uncertainty of the agent about its environment (Pathak et al., 2017; Burda et al., 2018; Zhang et al., 2021c), or that maximize the entropy of states and/or actions (Williams & Peng, 1991; Bellemare et al., 2016; Haarnoja et al., 2019; Hazan et al., 2019; Lee et al., 2019; Islam et al., 2019; Guo et al., 2021; Zhang et al., 2021b). Optimizing a surrogate objective is particularly effective for solving tasks with complex dynamics and reward functions, or with sparse rewards.
The differences between theory and practical implementations of exploration has led to common folklore seeking to provide intuition for the efficiency of policy-gradient methods. This work is part of the research line that studies the maximization of practical surrogate learning objective functions from a mathematical optimization perspective. Close to our work, studies of the learning objective with entropy regularization (an exploration-based reward shaping technique where the entropy of the policy is added in the learning objective) were conducted. It includes the study by Ahmed et al. (2019) concluding that it helps to provide smooth learning objective functions. The same exploration strategy was reinterpreted as a robust optimization method by Husain et al. (2021) and equivalently as a two-player game by Brekelmans et al. (2022). Bolland et al. (2023) furthermore argued that optimizing an entropy regularized objective is equivalent to optimizing the return of another policy with larger variance. More general studies on the learning dynamics have focused on the influence of baselines in policy gradient (Chung et al., 2021), and reward-shaping strategies that do not modify the learning objective, called potential based (Ng et al., 1999; Wiewiora et al., 2003; Harutyunyan et al., 2015; Forbes et al., 2024). All these studies are too restrictive and the literature lacks unified explanations and interpretations about exploration in policy gradients.
Before delving into our contributions, we recall that the convergence of stochastic ascent methods is driven by the objective function and how the ascent directions are estimated. First, the objective function shall be (pseudo) concave to find its global maximum (Bottou, 1998). Second, the convergence rate is influenced by the distribution of the stochastic ascent estimates (Chen & Luss, 2018; Ajalloeian & Stich, 2020). In this paper, we rigorously study policy-gradient methods with exploration-based reward shaping through the lens of these two optimization theory aspects. To that end, we first introduce two new criteria that relate the return of a policy to the learning objective with exploration bonuses, and their respective optima. Second, we introduce two additional criteria on the distribution of the gradient estimates of the learning objective and their likelihood of providing directions in which the learning objective and the return increase. Importantly, these criteria are general to any reward-shaping strategy, and highlight the importance of reward shaping that modify the optimal control behavior, in opposition to the literature on potential-based reward shaping. The influence of some exploration bonuses are illustrated and discussed in the light of these four criteria. In practice, finding good exploration strategies is problem specific and we thus introduce a general framework for the study and interpretation of exploration in policy-gradient methods instead of trying to find the best exploration method for a given task.
The paper is organized as follows. In Section 2, we provide the background about policy gradients and exploration. Section 3 focuses on the effect of exploration on the learning objective while Section 4 is dedicated to the effect on the gradient estimates used in the policy-gradient algorithms [^1] Experimental details and implementations can be found at
https://github.com/adrienBolland/micro-rl-lib.
. Conclusions and future works are discussed in Section 5.
2 Background
In this section, we introduce Markov decision processes and policy gradients with intrinsic exploration.
2.1 Markov Decision Processes
We study problems in which an agent makes sequential decisions in a stochastic environment (Sutton & Barto, 2018). The environment is modeled with an infinite-time Markov decision process (MDP) composed of a state space $\mathcal{S}$, an action space $\mathcal{A}$, an initial state distribution with density $p_{0}$, a transition distribution (modeling the dynamics) with conditional density $p$, a bounded reward function $\rho$, and a discount factor $\gamma\in[0,1)$. When an agent interacts with the MDP, first, an initial state $s_{0}\sim p_{0}(\cdot)$ is sampled, then, the agent provides at each time step $t$ an action $a_{t}\in\mathcal{A}$ leading to a new state $s_{t+1}\sim p(\cdot|s_{t},a_{t})$. Such a sequence of states and actions $h_{t}=(s_{0},a_{0},\dots,s_{t-1},a_{t-1},s_{t})\in\mathcal{H}$ is called a history and $\mathcal{H}$ is the set of all histories of any arbitrary length. In addition, after an action $a_{t}$ is executed, a reward $r_{t}=\rho(s_{t},a_{t})\in\mathbb{R}$ is observed.
A policy $\pi\in\Pi=\mathcal{S}\rightarrow\mathcal{P}(\mathcal{A})$ is a mapping from the state space $\mathcal{S}$ to the set of probability measures on the action space $\mathcal{P}(\mathcal{A})$, where $\pi(a|s)$ is the associated conditional probability density of action $a$ in state $s$. The function $J:\Pi\rightarrow\mathbb{R}$ is defined as the function mapping any policy $\pi$ to the expected discounted sum of rewards gathered by an agent interacting in the MDP by sampling actions from the policy $\pi$. We call return of the policy $\pi$ the value provided by that function
$J(\pi)=\frac{1}{1-\gamma}\underset{\begin{subarray}{c}s\sim d^{\pi,\gamma}(\cdot)\\ a\sim\pi(\cdot|s)\end{subarray}}{\mathbb{E}}\left[\rho(s,a)\right]\;,$ (1)
where $d^{\pi,\gamma}(\cdot)$ is the discounted state-visitation probability (Manne, 1960). In reinforcement learning, we seek to find an optimal policy $\pi^{*}$ maximizing the expected discounted sum of rewards $J$.
2.2 Policy-Gradient Algorithms
Policy-gradient algorithms (locally) optimize a parameterized policy $\pi_{\theta}$ to find the optimal parameter $\theta^{*}$ for which the return of the policy $J(\pi_{\theta^{*}})$ is maximized. Naively maximizing the return may provide sub-optimal results. This problem is mitigated in practice with exploration strategies, which consist in optimizing a surrogate learning objective $L$ that intrinsically encourages certain behaviors. In this work, we consider reward-shaping strategies where the expected discounted sum of rewards is extended by $K$ additional reward terms $\rho_{i}^{int}$, called intrinsic motivation terms, and optimize the learning objective
$L(\theta)$ $=\frac{1}{1-\gamma}\underset{\begin{subarray}{c}s\sim d^{\pi_{\theta},\gamma}(\cdot)\\ a\sim\pi_{\theta}(\cdot|s)\end{subarray}}{\mathbb{E}}\left[\rho(s,a)+\sum_{i=0}^{K-1}\lambda_{i}\rho_{i}^{int}(s,a)\right]$
$=J(\pi_{\theta})+J^{int}(\pi_{\theta})\;,$ (2)
where $\lambda_{i}$ are non-negative weights for each intrinsic reward and where $J^{int}(\pi_{\theta})$ is the intrinsic return of the policy. The parameter maximizing the learning objective is denoted by $\theta^{\dagger}$, which we distinguish from the optimal policy parameter $\theta^{*}$. Most of the intrinsic motivation terms can be classified in the two following groups.
Uncertainty-based motivations. It is common to provide bonuses for performing actions that reduce the uncertainty of the agent about its environment (Pathak et al., 2017; Burda et al., 2018; Zhang et al., 2021c). The intrinsic motivation terms are then proportional to the prediction errors of a model of the MDP dynamics. The latter model is usually learned.
Entropy-based motivations. It is also common to provide bonuses for visiting states and/or playing actions that are less likely in histories (Haarnoja et al., 2019; Hazan et al., 2019). In this work, we focus on two of these bonuses
$\rho^{s}(s,a)$ $=-\log d^{\pi_{\theta},\gamma}(\phi(s))$ (3)
$\rho^{a}(s,a)$ $=-\log\pi_{\theta}(a|s)\;,$ (4)
where $\phi(s)$ is a feature built from the state $s$. The corresponding intrinsic returns are maximized for policies that visit uniformly every feature, and for policies with uniformly distributed actions in each state, respectively. Note that these rewards require to estimate the distribution over the states and/or actions. Furthermore, they implicitly depend on the policy parameter $\theta$. The second technique is usually referred to as entropy regularization.
In this work, we consider on-policy policy-gradient algorithms, which were among others reviewed by (Duan et al., 2016) and (Andrychowicz et al., 2020). These algorithms optimize differentiable parameterized policies with gradient-based local optimization. They iteratively approximate an ascent direction $\hat{d}$ relying on histories sampled from the policy in the MDP and update the parameters in the ascent direction, or in a combination of the previous ascent directions (Hinton et al., 2012; Kingma & Ba, 2014). For the sake of simplicity and without loss of generality, we consider that the ascent direction $\hat{d}$ is composed of the sum of an estimate of the gradient of the return $\hat{g}\approx\nabla_{\theta}J(\pi_{\theta})$ and an estimate of the gradient of the intrinsic return $\hat{i}\approx\nabla_{\theta}J^{int}(\pi_{\theta})$. In practice, the first is usually unbiased while the second is computed neglecting some partial derivatives of $\theta$ and is thus biased, typically neglecting the influence of the policy on the intrinsic reward.
3 Study of the Learning Objective
In this section, we study the influence of the exploration terms on the learning objective defined in equation (2). We define two criteria under which the learning objective can be globally optimized by ascent methods, and such that the solution is close to an optimal policy. We then graphically illustrate how exploration modifies the learning objective to remove local extrema.
3.1 Policy-Gradient Learning Objective
Policy-gradient algorithms using exploration maximize the learning objective function $L$, as defined in equation (2). We introduce two criteria related to this learning objective for studying the performance of the policy-gradient algorithm. First, we say that a learning objective $L$ is $\epsilon$-coherent when its global maximum is in an $\epsilon$-neighborhood of the return of an optimal policy. Second, we call learning objectives that have a unique maximum and no other stationary point pseudoconcave.
Coherence criterion. A learning objective $L$ is $\epsilon$-coherent if, and only if,
$J(\pi_{\theta^{*}})-J(\pi_{\theta^{\dagger}})\leq\epsilon\;,$ (5)
where $\theta^{*}\in\text{argmax}_{\theta}J(\pi_{\theta})$ and where $\theta^{\dagger}\in\text{argmax}_{\theta}L(\theta)$.
Pseudoconcavity criterion. A learning objective $L$ is pseudoconcave if, and only if,
$\exists!\>\theta^{\dagger}:\nabla L(\theta^{\dagger})=0\land L(\theta^{\dagger})=\max_{\theta}L(\theta)\;.$ (6)
If the pseudoconcavity criterion is respected, there is a single optimum, and it is thus possible to globally optimize the learning objective function by (stochastic) gradient ascent (Bottou, 2010) [^2] For the sake of keeping discussions simple, the definition of pseudoconcavity is simplified (Mangasarian, 1975), and additional assumptions on the stochastic gradient estimates are neglected.. If the learning objective is furthermore $\epsilon$-coherent, the latter solution is also a near-optimal policy, where $\epsilon$ is the bound on the suboptimality of its return.
Let us finally remind a theorem from Ng et al. (1999).
Consistency Theorem. The learning objective $L$ is $\epsilon$-coherent, with $\epsilon=0$, in any MDP with state space $\mathcal{S}$, action space $\mathcal{A}$ and factor $\gamma$, if, and only if, $J(\theta)=L(\theta)$ for all $\theta$. The intrinsic rewards are furthermore potential based.
This theorem states that there is no MDP-agnostic exploration method that guarantees consistency with $\epsilon$ equal to zero and that modifies the objective function. This type of exploration is only possible with potential-based reward shaping (Ng et al., 1999). In conclusion, if the return is not pseudoconcave, there is a trade-off between the two criteria, which can not be resolved by potential-based exploration.
3.2 Illustration of the Effect of Exploration on the Learning Objective
Exploration is of paramount importance in environments with complex dynamics and reward functions, where many locally optimal policies may exist (Lee et al., 2019; Liu & Abbeel, 2021; Zhang et al., 2021b). In the following, we first define such an environment and a policy parameterization that will serve as an example to illustrate the effect of exploration on the optimization process. For the sake of the analysis, we then represent the learning objectives associated to different exploration strategies, and depict their global and local optima. Learning objectives with a single optimum respect the pseudoconcavity criterion. In addition, we represent the neighborhood $\Omega$ of the optimal policy parameters, such that any learning objective with its global maximum within this region is coherent for a given $\epsilon$. In light of the coherence and the pseudoconcavity criteria, we finally elaborate on the policy parameter computed by stochastic gradient ascent algorithms.
We consider the environment illustrated in Figure 1(a) where a car moves in a valley (Bolland et al., 2023). We denote by $x$ and $v$ the position and speed of the car, both composing its state $s=(x,v)$. The valley contains two separate low points, positioned in $x_{initial}=-3$ and $x_{target}=3$, separated by a peak. The car starts at rest $v_{0}=0$ at the highest low point $x_{0}=x_{initial}$ and receives rewards proportional to the depth of the valley at its current position. The reward function is provided in Figure 1(b). We consider a policy $\pi_{K,\sigma}(a|s)=\mathcal{N}(a|\mu_{K}(s),\sigma)$, namely a normally disturbed proportional controller with $\mu_{K}(s)=K\times(x-x_{target})$, parameterized by the vector $\theta=(K,\sigma)$. Figure 1(c) illustrates the contour map of the return of the policy as a function of the parameters $K$ and $\sigma$. The optimal parameters are represented by black dots and correspond to policies that drive the car to pass the peak and reach the lowest valley point in $x_{target}$. The green area represents the set of parameters $\Omega=\left\{\theta^{\prime}|\max_{\theta}J(\pi_{\theta})-J(\pi_{\theta^{\prime}})\leq\epsilon\right\}$ for $\epsilon=1$.
(a) Hill environment.
(b) Reward function $\rho$.
(c) Return $J(\pi_{K,\sigma})$.
Figure 1: Illustration of the hill environment in Figure 1(a) and its reward function in Figure 1(b). In Figure 1(c), the return of the policy $\pi_{K,\sigma}$ with the global and local maximum represented in black and grey, together with their respective return values.
Figure 2 illustrates the learning objective with the intrinsic rewards $\rho^{s}(s,a)=-\log d^{\pi_{K,\sigma},\gamma}(\phi(s))$, from equation (3), and $\rho^{a}(s,a)=-\log\pi_{K,\sigma}(a|s)$, from equation (4), for different values of the corresponding weights $\lambda_{s}$ and $\lambda_{a}$. Here, the feature is the position in the valley $\phi(s)=x$. First, we observe that for weights approaching zero, the parameter $\theta^{\dagger}$ maximizing the learning objective, represented by a black dot, corresponds to a policy with a high return. More precisely, it is in the green set $\Omega$ such that $\epsilon$-coherence is guaranteed for the value $\epsilon=1$. Larger weights require larger values of $\epsilon$ for guaranteeing the $\epsilon$-coherence criterion. Nevertheless, when increasing the weights, we observe that the learning objective eventually becomes pseudoconcave. There appears to be a trade-off between the two criteria. In Figure 2(b), we observe that in this environment, there is a learning objective that respects the pseudoconcavity criterion and the $\epsilon$-coherence criterion for $\epsilon=1$. Indeed, there is a single global maximum in Figure 2(b) represented by a black dot that is furthermore part of the set $\Omega$.
Shaping the reward function with an exploration strategy based on the state-visitation entropy appears to be a good solution for optimizing the policy. However, a notable drawback is that the reward depends on the policy and its (gradient) computation requires to estimate a complex probability measure. In this example, the intrinsic reward function itself was estimated by Monte-Carlo sampling for every parameter, which would not scale for complex problems and requires approximations and costly evaluation strategies (Islam et al., 2019). In Appendix A we present an alternative problem-dependent intrinsic reward, independent of the policy parameters and thus simple to compute efficiently, that still respects the pseudoconcavity and $\epsilon$-coherence criteria, and in Appendix B we extend the study to more complex environments from the MiniGrid library (Chevalier-Boisvert et al., 2023) where the policy is a deep neural network and the state-visitation probability is approximated.
(a) $\lambda_{s}=0.05$ and $\lambda_{a}=0$
(b) $\lambda_{s}=0.1$ and $\lambda_{a}=0$
(c) $\lambda_{s}=1$ and $\lambda_{a}=0$
(d) $\lambda_{s}=0$ and $\lambda_{a}=0.01$
(e) $\lambda_{s}=0$ and $\lambda_{a}=0.1$
(f) $\lambda_{s}=0$ and $\lambda_{a}=0.5$
Figure 2: Contour map of (scaled) learning objective functions for different values of $\lambda_{s}$ and $\lambda_{a}$. The darker the map, the larger the learning objective value. The green area represents the set $\Omega=\left\{\theta^{\prime}|\max_{\theta}J(\pi_{\theta})-J(\pi_{\theta^{\prime}})\leq\epsilon=1\right\}$, such that when the parameter maximizing the learning objective is part of $\Omega$, then the learning objective function is $\epsilon$-coherent with $\epsilon=1$. The black dot is the parameter $\theta^{\dagger}$ globally maximizing the learning objective and the grey dot is the local (non-global) maximum of the learning objective if it exists. Both are labeled with the return values of the corresponding policies.
The observations suggest that well-chosen exploration strategies can lead to learning objective functions that satisfy the two criteria defined in the previous section, thereby guaranteeing that policies suboptimal by at most $\epsilon$ can be computed by local optimization. When designing exploration strategies, it is essential to keep in mind that we modify the learning objective for the algorithms to converge to optimal policy parameters, which can be achieved when both criteria are respected. While strategies such as enforcing entropy can be effective in some environments, they are only heuristic strategies and not to be relied upon exclusively. Furthermore, as illustrated, both criteria may be subject to a trade-off. In more complex environments, an efficient exploration strategy may require to balance both criteria, e.g., through a schedule on the learning objective weights.
4 Study of the Ascent Direction Distribution
Optimizing pseudoconcave functions with stochastic ascent methods are guaranteed to converge (at a certain rate) under assumptions on the distribution of the gradient estimates at hand (Bottou, 2010; Chen & Luss, 2018; Ajalloeian & Stich, 2020). In this section, we study the influence of the exploration terms on this distribution in the context of policy gradients. More precisely, we study the probability of improving the learning objective and the return with stochastic ascent steps. Intuitively, they shall be sufficiently large for the algorithm to be efficient. We formalize this intuition and illustrate how exploration strategies can increase these probabilities, leading to more efficient algorithms.
4.1 Policy-Gradient Estimated Ascent Direction
In general, gradient ascent algorithms update parameters in a direction $\hat{d}$ in order to locally improve an objective function $f$. The quality of these algorithms can therefore be studied (for a small step size $\alpha\rightarrow 0$) through the random variable representing the quantity by which the objective increases for each $\theta$
$X$ $=f(\theta+\alpha\hat{d})-f(\theta)=\alpha\>\langle\hat{d},\nabla_{\theta}f(\theta)\rangle\>,$ (7)
[…]
The (asymptotic) convergence of a gradient ascent algorithm is usually studied by bounding the expectation of $X$. Such bounds depend, among others, on the expectation of $\hat{d}$, which equals $\nabla_{\theta}f(\theta)$ when unbiased, and depend on the variance of $\hat{d}$, which deteriorates the expected convergence rate. Finding rates depending on the algorithm is an active field of research. In parallel, a slightly more general problem is to quantify if the expectation of $X$ is driven by rare events. Intuitively, an algorithm with unbiased gradient estimates has positive expected improvements $\mathbb{E}[X]>0$, and should theoretically converge, but may be inefficient in practice if positive events $X>0$ rarely occur. We illustrate in the next section that this phenomenon makes reinforcement learning in sparse-reward environments particularly hard. To the best of our knowledge, no existing result fits to the study of policy gradients. We therefore introduce two new criteria on the probability of improvement $P(X>0)$, which we empirically validate afterwards.
First, we define an exploration strategy as $\delta$-efficient if, and only if, following the ascent direction $\hat{d}\approx\nabla_{\theta}L(\theta)$ has a probability at least $\delta$ to increasing the learning objective $L(\theta)$ almost everywhere. Second, an exploration strategy is $\delta$-attractive if, and only if, there exists a neighborhood of $\theta^{\dagger}$ containing the parameter $\theta^{int}$ maximizing the intrinsic return $J^{int}$, where the probability of increasing the return by following $\hat{d}$ is almost everywhere at least equal to $\delta$. Note that each probability measure and random variable is a function of $\theta$, which we do not explicitly write for the sake of keeping notations simple.
Efficiency criterion. An exploration strategy is $\delta$-efficient if, and only if,
$\forall^{\infty}\theta:\mathbb{P}(D>0)\geq\delta\>,$ (8)
where $D=\>\langle\hat{d},\nabla_{\theta}L(\theta)\rangle$.
Attraction criterion. An exploration strategy is $\delta$-attractive if, and only if,
$\exists B(\theta^{\dagger})$ $:\theta^{int}\in B(\theta^{\dagger})\>,$ (9)
such that
$\forall^{\infty}\theta\in B(\theta^{\dagger}):\mathbb{P}(G>0)\geq\delta\>,$ (10)
where $\theta^{int}=\text{argmax}_{\theta}J^{int}(\pi_{\theta})$, $B(\theta^{\dagger})$ is a ball centered in $\theta^{\dagger}$, and $G=\>\langle\hat{d},\nabla_{\theta}J(\pi_{\theta})\rangle$.
The efficiency criterion quantifies how often a stochastic gradient ascent step improves the learning objective. The larger, the better the learning objective and its stochastic ascent direction approximations. The rationale behind the attraction criterion is that in many exploration strategies, the intrinsic reward is dense, and it is then presumably easy to optimize the intrinsic return in the sense that $\mathbb{P}(\>\langle\hat{i},\nabla_{\theta}J^{int}(\pi_{\theta})\rangle>0)$ is large. It implies that it is easy to locally improve the learning objective by (solely) increasing the value of the intrinsic motivation terms. It furthermore implies that policy-gradient algorithms may be subject to converging towards $\theta^{int}$ rather than $\theta^{\dagger}$ when $\mathbb{P}(\>\langle\hat{d},\nabla_{\theta}J(\pi_{\theta})\rangle>0)$ is small. If the criterion is respected for large $\delta$, the latter is less likely to happen as policy gradients will eventually tend to improve the return of the policy if the parameter approaches $\theta^{int}$ and enters the ball $B(\theta^{\dagger})$; eventually converging towards $\theta^{\dagger}$.
These two new criteria on $\hat{d}$ are independent of the previous ones on $L$, which only captured the quality of the deterministic learning objective functions. In the particular cases where the learning objectives $L$ are $\epsilon$-coherent, for $\epsilon=0$, and pseudoconcave, e.g., with potential-based intrinsic rewards, only the distribution of estimates $\hat{d}$ can explain why some algorithms succeed and others fail. Finally, the value of $\delta$ in the new criteria we introduce can be related to the variance of the estimate $\hat{d}$ under some assumptions, e.g., with Cantelli’s concentration inequalities.
4.2 Illustration of the Effect of Exploration on the Estimated Ascent Direction
Exploration is usually promoted and tested for problems where the reward function is sparse, typically in maze-environments (Islam et al., 2019; Liu & Abbeel, 2021; Guo et al., 2021). In this section, we first introduce a new maze-environment with sparse rewards where we illustrate the influence of exploration on the gradient estimates of the learning objective. To this end, we present two learning objective functions and elaborate on the influence of exploration on the performance of policy-gradient algorithms in the light of the efficiency and attraction criteria.
Let us consider a maze-environment consisting of a horizontal corridor composed of $S\in\mathbb{N}$ tiles. The state of the environment is the index of the tile $s\in\left\{1,\dots,S\right\}$, and the actions consist in going left $a=-1$ or right $a=+1$. When an action is taken, the agent stays idle with probability $p=0.7$, and moves with probability $0.31-p=$ 0.3 in the direction indicated by the action, then $s^{\prime}=\min(S,\max(1,s+a))$. The agent starts in state $s=1$ and the target state $s=S=15$ is absorbing. Zero rewards are observed except when the agent reaches the target state where a reward $r=100$ is observed. A discount factor of $\gamma=0.99$ is considered. Finally, we study the policy going with probability $\theta$ to the right and probability $1-\theta$ to the left, and with density
$\pi_{\theta}(a|s)=\left\{\begin{array}[]{cl}\theta&\quad\textrm{if }a=1\\ 1-\theta&\quad\textrm{if }a=-1\;.\end{array}\right.$ (13)
The return $J(\pi_{\theta})$ is represented in black in Figure 3(a) as a function of $\theta$ along with two intrinsic returns, $J^{a}(\pi_{\theta})$ in orange and $J^{s}(\pi_{\theta})$ in blue. The intrinsic reward $\rho^{a}(s,a)=-\log\pi_{\theta}(a|s)$, from equation (4), and the intrinsic reward $\rho^{s}(s,a)=-\log d^{\pi_{\theta},\gamma}(s)$, from equation (3), are used respectively. In Figure 3(b), we illustrate the return of the policy without exploration $J(\pi_{\theta})$, along with two learning objective functions, $L^{a}(\theta)$ and $L^{s}(\theta)$, using as exploration strategies the intrinsic returns $J^{a}(\pi_{\theta})$ and $J^{s}(\pi_{\theta})$. We observe that the return is a pseudoconcave function with respect to $\theta$ and the optimal parameter is $\theta^{*}=1$. In addition, the two learning objectives respect the $\epsilon$-coherence criterion for $\epsilon=0$, implying that $\theta^{*}=\theta^{\dagger}$, and respect the pseudoconcavity criterion. It is important to note that with regard to the discussion from Section 3, there is no interest in optimizing the learning objectives rather than directly optimizing the return, as the latter is already pseudoconcave. In the following we illustrate how choosing a correct exploration strategy still deeply influences the policy-gradient algorithms when it comes to building gradient estimates.
(a) Return functions.
(b) Learning objective functions.
(c) Improvement probabilities.
Figure 3: Figure 3(a) represents the return of the policy along with two intrinsic return functions. In Figure 3(b) the return is also represented together with two learning objective functions, corresponding to the two intrinsic returns. Figure 3(c) illustrates the probability (estimated by Monte-Carlo) of positive stochastic gradient (derivative) estimates $J(\pi_{\theta})$, $L^{a}(\theta)$, and $L^{s}(\theta)$. At the top of the figure, the intervals $B^{a}=[\theta^{int,a},\theta^{\dagger,a}]$ and $B^{s}=[\theta^{int,s},\theta^{\dagger,s}]$ are represented. These intervals represent the smallest balls containing the parameters maximizing the intrinsic return and the learning objective, for both exploration strategies.
Let us compute the estimate $\hat{g}$ and $\hat{d}$ relying on REINFORCE (Williams, 1992) by sampling $8$ histories of length $T=100$. In this particular environment, $\mathbb{P}(D>0)$ equals $\mathbb{P}(G>0)$, and equal the probability that the derivative is positive. We represent in Figure 3(c) this probability for the return and for both learning objectives. First, we see that the learning objectives are more efficient than the return, meaning they are $\delta$-efficient for larger values of $\delta$. Depending on the parameter value, the objective $L^{a}(\theta)$ or $L^{s}(\theta)$ is best in that regard. Second, concerning the attraction criterion, we represent at the top of Figure 3(c) the intervals $B^{a}=[\theta^{int,a},\theta^{\dagger,a}]$ and $B^{s}=[\theta^{int,s},\theta^{\dagger,s}]$. They correspond to the smallest balls containing the maximizers of the intrinsic return and of the learning objective. Let the minima of the orange and blue curves over these intervals be denoted by $\delta^{a}$ and $\delta^{s}$. By definition of the attraction criterion, it is thus respected for any values of $\delta$ at most equal to $\delta^{a}$ and $\delta^{s}$, for $L^{a}(\theta)$ and $L^{s}(\theta)$, respectively. All these observations can eventually be explained as the computation of $\hat{g}$ is always zero when the target is not sampled in the histories, which is highly likely for policies with small values of $\theta$. Policy-gradient algorithms relying on intrinsic exploration would compute optimal policies efficiently where naive optimization without exploration would fail or be sample inefficient.
We have empirically shown that a well-chosen exploration strategy in policy gradients may not only remove local extrema from the objective function, but may also increase the probability that stochastic ascent steps improve the objective function. Under the previous assumptions, this probability measures the efficiency of algorithms. Furthermore, among different learning objectives respecting the coherence and pseudoconcavity criteria, it is best to choose one that has high values for $\delta$ in both the efficiency and attraction criteria. In Appendix A we use these criteria to study other reward-shaping strategies, and in Appendix B we extend the study to more complex environments from the MiniGrid library (Chevalier-Boisvert et al., 2023) where the policy is a deep neural network.
The problem discussed in this section strongly relates to overfitting or generalization in reinforcement learning. In situations where the same state and action pairs are repeatedly sampled with high probability, the policy may appear optimal by neglecting the rewards observed in state and action pairs sampled with low probability. The gradient estimates will then be zero with high probability, and the gradient updates will not lead to policy improvements. In the previous example, gradient estimates computed from policies with a small parameter value $\theta$ wrongly indicate that a stationary point has been reached as they equal zero with high probability. We quantify this effect with a novel definition of local optimality. We define as locally optimal policies over a space with probability $\Delta$ the policies that maximize the reward on expectation over a set of states and actions observed in a history with probability at least $\Delta$. Formally, a policy $\pi$ is locally optimal over a space with probability $\Delta$ if, and only if,
[…]
In the typical case of environments with sparse rewards, many policies observe with high probability state and action pairs with zero rewards and are locally optimal for large probabilities $\Delta$. Typically, in the previous example, the joint set $\{1,\dots,S-2\}\times\{-1,1\}$ is a set of state and action pairs $\mathcal{E}$ that respects the definition equation (14) for large values $\Delta$ when $\theta$ is small. As we have shown, exploration mitigates the convergence of policy-gradient algorithms towards these locally optimal policies. Note that assuming a non-zero reward is uniformly distributed over the state and action space, exploration policies with uniform probabilities over visited states and actions are the best choice for sampling non-zero rewards with high probability. It can thus also be considered as the best choice of exploration to reduce the probability that the stochastic gradient ascent steps do not increase the objective value. Such initial policy may be learned from the framework developed by Lee et al. (2019).
5 Conclusion
In conclusion, this research takes a step towards dispelling misunderstandings about exploration through the study of its effects on the performance of policy-gradient algorithms. More particularly, we distinguished two effects exploration has on the optimization. First, it modifies the learning objective in order to remove local extrema. Second, it modifies the gradient estimates and increases the likelihood that the update steps lead to improved returns. These two phenomena were studied through four criteria that we introduced and illustrated.
These ideas apply to other direct policy optimization algorithms. Indeed, the four criteria do not assume any structure on the learning objective and can thus be straightforwardly applied to any objective function optimized by a direct policy search algorithm. In particular, for off-policy policy gradient, we may simply consider that the off-policy objective is itself a surrogate or that the gradients of the return are biased estimates based on past histories. Ideas introduced in this work also apply to other reinforcement learning techniques. Typically, for value-based RL with sparse-reward environments, convergence towards a value function outputting zero is expected with high probability. This is mostly due to the low probability of sampling non-zero rewards by Monte-Carlo. The discussions from Section 4 then apply, and a similar analysis can be performed.
Our framework opens the door for further theoretical analysis, and the potential development of new criteria. We believe that deriving practical conditions on the exploration strategies, and the scheduling of the intrinsic return, for guaranteeing fast convergence should be the focus of attention. It could be achieved by bounding the policy improvement on expectation, which is nevertheless usually a hard task without strong assumptions. We furthermore believe that we provide a new lens on exploration necessary for interpreting and developing exploration strategies, in the sense of optimizing surrogate learning objective functions.
Acknowledgments
The authors thank Arnaud Delaunoy, Pascal Leroy, and Mathias Berger for valuable comments on this manuscript. Adrien Bolland and Gaspard Lambrechts gratefully acknowledge the financial support of the F.R.S.-FNRS.
Appendix A Reward Shaping and Exploration Strategies
As discussed in the manuscript, exploration strategies are reward-shaping strategies where the intrinsic reward bonuses are, among others, dependent on the policy parameters. This dependency makes the shaping strategies adaptive but makes the computation of gradients and the study of the learning objectives more complex. In this section, we study handcrafted reward-shaping strategies that have pseudoconcave and dense reward functions in the hill and maze environments. We then illustrate that the same criteria can be used to study these expert-knowledge based shaped rewards.
For the hill environment from Section 3, we illustrate in Figure 4(a) an intrinsic reward bonus making the sum of rewards in equation (2) concave. The corresponding learning objective has a unique maximum, which is part of the set $\Omega=\left\{\theta^{\prime}|\max_{\theta}J(\pi_{\theta})-J(\pi_{\theta^{\prime}})\leq\epsilon\right\}$ with $\epsilon=1$ and $\theta=(K,\sigma)$. It can be seen in Figure 4(b) where the global maximum in black is within the set $\Omega$ in green. Both, the $\epsilon$-coherence and the pseudoconcavity criteria are thus respected for $\epsilon=1$. Here, the intrinsic reward function is a simple function independent of the policy $\pi_{\theta}$. Finding such an intrinsic reward may be complex for other environments but the example underlines that exploration and reward shaping are mostly equivalent and that designing reward functions that are concave may help converging towards optimal policies.
(a) Alternative reward function.
(b) Learning objective.
Figure 4: In Figure 4(a), an alternative intrinsic reward function ensuring that the sum of rewards is a pseudoconcave function. In Figure 4(b), the contour function of the learning objective.
For the maze environment, the return $J(\pi_{\theta})$ is represented in black in Figure 5(a) together with the intrinsic return $J^{d}(\pi_{\theta})$ in green. The latter is the return of the dense handcrafted reward function $\rho^{d}(s,a)=(a-1)/2$ penalizing actions moving away from the target. In Figure 5(b), the corresponding learning objective function is shown. In the same experimental setting as in Section 4, we observe that the objective function is $\delta$-efficient for higher values of $\delta$ compared to the already-discussed learning objectives. Furthermore, the attraction criterion is respected for any value of $\delta$ as the unique global maxima of the learning objective, intrinsic return, and return are all equals.
(a) Return.
(b) Learning objectives.
(c) Improvement probabilities.
Figure 5: In Figure 5(a) the return of the maze environment is represented together with the intrinsic return of a dense handcrafted reward function. Figure 5(b) represents the corresponding learning objective and Figure 5(c) the probability that the REINFORCE estimates are positive.
Appendix B Minigrid Experiments
In this section, we introduce complex environments and parameterize policies with neural networks. In this context, it is impractical to naively compute and represent the objective functions and probability distributions for the different criteria. Therefore, we only evaluate the criteria along parameter trajectories, and extend the previous experimental setting.
We consider seven environments from the MiniGrid suite of environments (Chevalier-Boisvert et al., 2023), among others, designed for evaluating exploration strategies. In these environments, an agent moves in a maze and aims to reach a target position. To do so, the agent may choose actions that consist of turning left, turning right, moving forward, or staying idle. We consider two reward settings: the dense setting and the sparse setting. In the first, rewards of $-1$ are received for every non-idle move, and a reward of $1000$ is received upon reaching the target position. In the second setting, zero rewards are received everywhere, except upon reaching the target position, where a bonus of $1000$ is provided. In the dense setting, due to the action penalization incurred when moving, a policy outputting the idle action with probability one is locally optimal and has a return equal to zero. This is not (necessarily) the case in the second setting. We consider a discount factor of $\gamma=0.98$ and optimize a fully connected neural network taking as input the position pair and the orientation of the agent, and outputting a categorical distribution over actions. The network is composed of three hidden layers of $64$ neurons with ReLU activation functions.
In the dense reward setting, we optimize policies by maximizing three learning objective functions: $J(\pi_{\theta})$, $L^{a}(\theta)$, and $L^{s}(\theta)$, respectively with $\lambda_{a}=0.5$ and $\lambda_{s}=0.25$. For the last objective, the state-visitation density estimator is a ten-component Gaussian mixture model maximizing the likelihood of the sampled batch. The optimization is performed using the Adam update rule (Kingma & Ba, 2014), with REINFORCE ascent directions computed over 32 histories of constant length $T=100$, and with learning rate (step size) equal to $0.0005$. The length $T$ of the histories is chosen such that the realization value $T$ from a geometric distribution with success probability parameter $1-\gamma$ has at least a cumulative probability of $0.85$. In this setting, we illustrate the quasiconcavity criterion and the $\epsilon$-coherence criterion. In Figure 6, we provide the evolution of the return of the policies when optimizing the three objectives $J(\pi_{\theta})$, $L^{a}(\theta)$, and $L^{s}(\theta)$ for the different environments. For the MiniGrid-Empty-8x8-v0 and the MiniGrid-FourRooms-v0 environments, optimizing the return results in high-performance policies that do not stay idle. The other objectives also manage to find high-performing policies, but with a lower return. This phenomenon (assuming that the global optimum of each objective is found) illustrates the $\epsilon$-coherence criterion, where this $\epsilon$ value is the bound on the best policy that can be found when optimizing the learning objective. For the other environments, the policies resulting from the optimization of the return fall into local optima, namely ones where the policy chooses the idling action with probability one. When optimizing the learning objectives with exploration bonuses, the resulting policies no longer fall into the previous local optima. This result suggests that, along these parameter trajectories, the return $J(\pi_{\theta})$ has a local optimum (or saddle point), in opposition to the learning objective functions $L^{a}(\theta)$ and $L^{s}(\theta)$. The latter illustrates the validity of the pseudoconcavity criterion in that region of the parameter space. For the learning objective $L^{s}(\theta)$, the $\epsilon$-coherence criterion is respected for a small value of $\epsilon$, and the resulting policy manages to reach the target position. For the learning objective $L^{a}(\theta)$, the resulting policy does not reach the target position. We nevertheless hypothesize that it is not a real local optimum of the learning objective but a local optimum in the sense of equation (14).
Figure 6: Evolution of the return of policies during optimization in the dense minigrid environments. In blue, the return $J(\pi_{\theta})$ is optimized; in orange, the learning objective $L^{a}(\theta)$ is optimized; and in green, the learning objective $L^{s}(\theta)$ is optimized by performing Adam steps in REINFORCE directions. Note that the median, worst, and best cases over five runs are represented for the different curves. For the environments MiniGrid-Empty-8x8-v0 and MiniGrid-FourRooms-v0, optimizing each objective results in a policy that does not stay idle. These policies are initialized outside of the basin of attraction of the local optimum of the return. The coherence criteria can be observed as optimizing the learning objective with intrinsic exploration bonuses results in suboptimal policies. For the other environments, optimizing the return directly leads to policies that always choose to stay idle, and are thus locally optimal. Optimizing the learning objective with exploration allows us to escape from these local optima, illustrating the quasiconcavity criterion. It can be noted that when optimizing the objective $L^{a}(\theta)$, the $\epsilon$-coherence criterion is respected for a large value of $\epsilon$, making the resulting policy worse than when optimizing the return directly. On the contrary, the objective $L^{s}(\theta)$ appears to have $\epsilon$-coherence for a reasonable value of $\epsilon$.
In the previous experiments with the dense setting, the local optima exist due to the negative rewards associated to idle-actions. If we consider the sparse setting, we could then assume that directly optimizing the return is sufficient to find high-performing policies. It is not always the case and it can be justified using the efficiency and attraction criteria. We use the same parameters as in the previous set of experiments and we provide the evolution of the return of the policies when optimizing the three objectives $J(\pi_{\theta})$, $L^{a}(\theta)$, and $L^{s}(\theta)$, for the different environments in Figure 7. On the one hand, for most environments, whatever the learning objective, the resulting policy has a high return. Note that the $\epsilon$-coherence can again be illustrated where the policies resulting from the optimization of the return perform better than the others. On the other hand, for the MiniGrid-Empty-16x16-v0 and the MiniGrid-SimpleCrossingS11N5-v0 environments, a high performing policy can only be found when optimizing the learning objective $L^{s}(\theta)$. We illustrate that these results can be justified by the efficiency and attraction criteria in Figure 8. For each parameters obtained during the stochastic ascent steps on the return, we first estimate the probability of improving both objective functions by stochastic gradient ascent. These probabilities are used to compare learning objectives in terms of the efficiency criterion; the larger, the better. Then, in order to illustrate the attraction criterion, we estimate the probability of improving the return and the learning objective, both by gradient ascent steps on the learning objective, for each parameters obtained during the stochastic ascent steps on the objective $L^{a}(\theta)$ and $L^{s}(\theta)$. As far as $L^{a}(\theta)$ is concerned, all probabilities remain small as result that the optimization procedure converges fast towards a stationary point where the target goal is observed with negligible probability. The efficiency and attraction criteria are respected for negligible probabilities $\delta$, which is also a justification for the failure of converging towards good policies in the dense setting. For $L^{s}(\theta)$, on the contrary, and in both environments, the probability of improving the learning objective remains large for each parameter encountered when optimizing the return. The efficiency of the learning objective is much higher than that of the return in that part of the parameter space. Furthermore, the probability of improving the return when optimizing the learning objective, is small at the beginning and increases after some iterations. This indicates that once the policy has a sufficiently large intrinsic return, the attraction criterion is respected for a high value $\delta$.
Figure 7: Evolution of the return of policies during optimization in the sparse minigrid environments. In blue, the return $J(\pi_{\theta})$ is optimized; in orange, the learning objective $L^{a}(\theta)$ is optimized; and in green, the learning objective $L^{s}(\theta)$ is optimized performing Adam steps in REINFORCE directions. Note that the median, worst, and best cases over five runs are represented for the different curves. In most environments, a high-performing policy can be found by optimizing the return $J(\pi_{\theta})$. This results from its quasiconcavity. The $\epsilon$-coherence criterion can also be observed. However, for the MiniGrid-Empty-16x16-v0 and the MiniGrid-SimpleCrossingS11N5-v0 environments, a high-performing policy can only be found when optimizing the learning objective $L^{s}(\theta)$.
Figure 8: For the MiniGrid-Empty-16x16-v0 and the MiniGrid-SimpleCrossingS11N5-v0 environments, and for both learning objective functions $L^{a}(\theta)$ and $L^{s}(\theta)$, we first represent the estimated probability of improving the return $J(\pi_{\theta})$ and the corresponding learning objective when following their REINFORCE gradient estimate. This value is estimated at each run of the optimization of the policy with learning objective $J(\pi_{\theta})$. Second, we represent the estimated probability of improving the corresponding learning objective and the return $J(\pi_{\theta})$ when following the REINFORCE gradient estimate of the learning objectives. These values are estimated at each run of the optimization of the policy with the learning objectives. The probabilities were estimated based on the frequencies of improving the objective functions by more than $0.2$ when following $5$ Adam ascent steps using REINFORCE update directions.

### v3

@@
 Abstract
-Policy-gradient algorithms are effective reinforcement learning methods for solving control problems.
-To compute near-optimal policies, it is essential in practice to include exploration terms in the learning objective.
+In order to compute near-optimal policies with policy-gradient algorithms, it is common in practice to include intrinsic exploration terms in the learning objective.
 Although the effectiveness of these terms is usually justified by an intrinsic need to explore environments, we propose a novel analysis with the lens of numerical optimization.
 Two criteria are introduced on the learning objective and two others on its stochastic gradient estimates, and are afterwards used to discuss the quality of the policy after optimization.
-The analysis sheds the light on two separate effects of exploration techniques.
+The analysis sheds light on two separate effects of exploration techniques.
 First, they make it possible to smooth the learning objective and to eliminate local optima while preserving the global maximum.
 Second, they modify the gradient estimates, increasing the probability that the stochastic parameter updates eventually provide an optimal policy.
-These effects are illustrated empirically on exploration strategies based on entropy bonuses, highlighting their limitations and opening avenues for future works in the design and analysis of such strategies.
-Reinforcement Learning, Policy Gradients, Exploration
+We empirically illustrate these effects with exploration strategies based on entropy bonuses, identifying limitations and suggesting directions for future work.
 1 Introduction
-Many practical problems require making sequential decisions in environments.
-Reinforcement learning (RL) is a framework for solving such decision-making problems that has been successful on complex tasks, including playing games (Mnih et al., 2015; Silver et al., 2017), operating power systems (Aittahar et al., 2024), controlling robots (Kalashnikov et al., 2018), or interacting with electricity markets (Boukas et al., 2021).
+Reinforcement learning (RL) has been successful in complex decision-making problems, including playing games (Mnih et al.,, 2015; Silver et al.,, 2017), operating power systems (Aittahar et al.,, 2024), controlling robots (Kalashnikov et al.,, 2018), and trading electricity (Boukas et al.,, 2021).
 Reinforcement learning algorithms interact with an environment to gather information about this environment, which in turn enables to compute and follow an optimal policy.
-This creates a trade-off between exploration and exploitation.
-In short, in order to eventually compute a good policy, it is necessary to obtain additional information about the environment by taking actions that are likely not optimal.
-In algorithms where the trade-off is explicit, exploration is well-understood and has been the subject of many works (Dann et al., 2017; Azar et al., 2017; Neu & Pike-Burke, 2020).
-In policy-gradient algorithms, one can most often not distinguish exploration from exploitation.
-Nevertheless, a main theoretical requirement to converge towards globally (or even locally) optimal solutions is that policies remain sufficiently stochastic during the learning procedure (Bhandari & Russo, 2019; Bhatt et al., 2019; Agarwal et al., 2020; Zhang et al., 2021a; Bedi et al., 2022).
-Interestingly, neither softmax nor Gaussian policies guarantee enough stochasticity for ensuring (fast) convergence (Mei et al., 2020, 2021; Bedi et al., 2022).
-This requirement of stochasticity in policy-gradient algorithms is often abusively called exploration and understood as the need to infinitely sample all states and actions.
-Practitioners have tried to meet the theoretical requirement of sufficient randomness of policies in policy gradient via reward-shaping strategies, whereby a learning objective that promotes or hinders behaviors by providing reward bonuses for some states and actions is optimized as a surrogate to the return of the policy.
-These bonuses typically promote actions that reduce the uncertainty of the agent about its environment (Pathak et al., 2017; Burda et al., 2018; Zhang et al., 2021c), or that maximize the entropy of states and/or actions (Williams & Peng, 1991; Bellemare et al., 2016; Haarnoja et al., 2019; Hazan et al., 2019; Lee et al., 2019; Islam et al., 2019; Guo et al., 2021; Zhang et al., 2021b).
-Optimizing a surrogate objective is particularly effective for solving tasks with complex dynamics and reward functions, or with sparse rewards.
-The differences between theory and practical implementations of exploration has led to common folklore seeking to provide intuition for the efficiency of policy-gradient methods.
-This work is part of the research line that studies the maximization of practical surrogate learning objective functions from a mathematical optimization perspective.
-Close to our work, studies of the learning objective with entropy regularization (an exploration-based reward shaping technique where the entropy of the policy is added in the learning objective) were conducted.
-It includes the study by Ahmed et al.
-(2019) concluding that it helps to provide smooth learning objective functions.
-The same exploration strategy was reinterpreted as a robust optimization method by Husain et al.
-(2021) and equivalently as a two-player game by Brekelmans et al.
-(2022).
-Bolland et al.
-(2023) furthermore argued that optimizing an entropy regularized objective is equivalent to optimizing the return of another policy with larger variance.
-More general studies on the learning dynamics have focused on the influence of baselines in policy gradient (Chung et al., 2021), and reward-shaping strategies that do not modify the learning objective, called potential based (Ng et al., 1999; Wiewiora et al., 2003; Harutyunyan et al., 2015; Forbes et al., 2024).
-All these studies are too restrictive and the literature lacks unified explanations and interpretations about exploration in policy gradients.
+This interaction creates a trade-off between exploration and exploitation.
+In short, to eventually compute a good policy, an agent must obtain additional information about the environment by taking actions that are likely not optimal.
+In algorithms where this trade-off is explicit, exploration is well understood and has been the subject of many works (Dann et al.,, 2017; Azar et al.,, 2017; Neu and Pike-Burke,, 2020).
+In policy-gradient algorithms, one often can not distinguish exploration from exploitation.
+Nevertheless, a key theoretical requirement for convergence to globally (or even locally) optimal solutions is that policies remain sufficiently stochastic during the learning procedure (Bhandari and Russo,, 2019; Bhatt et al.,, 2019; Agarwal et al.,, 2020; Zhang et al., 2021a, ; Bedi et al.,, 2022).
+Interestingly, neither softmax nor Gaussian policies guarantee enough stochasticity to ensure (fast) convergence (Mei et al.,, 2020, 2021; Bedi et al.,, 2022).
+This requirement for stochasticity in policy gradients is often loosely called exploration and understood as the need to infinitely sample each state-action pair.
+The need for sufficient policy randomness has long been known by practitioners, who have attempted to meet this requirement in policy-gradient methods through reward-shaping strategies.
+In these approaches, a learning objective that promotes or hinders behaviors by assigning reward bonuses to some states and actions is optimized as a surrogate for the policy return.
+Such bonuses typically encourage actions that reduce uncertainty of the agent about its environment (Pathak et al.,, 2017; Burda et al.,, 2018; Zhang et al., 2021c, ) or that maximize the entropy of states or actions (Williams and Peng,, 1991; Bellemare et al.,, 2016; Haarnoja et al.,, 2019; Hazan et al.,, 2019; Lee et al.,, 2019).
+Optimizing surrogate learning objectives that include reward bonuses appears particularly effective for solving challenging tasks, typically those with complex dynamics or sparse rewards.
+The connection between such objectives and exploration has nevertheless long been limited to intuition, and the lack of theoretical analysis has led to common folklore that seeks to justify their practical efficiency.
+Research on these objectives has mainly focused on entropy regularization (an exploration-based reward-shaping technique in which the policy entropy is added to the learning objective).
+This line of work includes the empirical study of Ahmed et al., (2019), which concludes that entropy regularization helps to provide smoother objective functions.
+The same strategy has been reinterpreted as a robust optimization method by Husain et al., (2021) and, equivalently, as a two-player game by Brekelmans et al., (2022).
+Bolland et al., (2023) further argue that optimizing an entropy-regularized objective is equivalent to maximizing the return of another policy with larger variance.
+Finally, convergence rates for policy-gradient algorithms with entropy regularization have recently been derived (Cen et al.,, 2022; Zhang et al., 2021a, ).
+More general studies of learning dynamics have examined the influence of baselines in policy-gradient methods (Chung et al.,, 2021) and potential-based reward-shaping strategies that do not modify the learning objective (Ng et al.,, 1999; Wiewiora et al.,, 2003; Harutyunyan et al.,, 2015; Forbes et al.,, 2024).
+All of these studies remain restricted to particular learning objectives and the literature still lacks unified explanations of the improvements intrinsic reward bonuses can provide, the conditions under which they arise, and systematic ways to compare different bonuses, among others.
+This work aims to take a step toward answering these questions.
 Before delving into our contributions, we recall that the convergence of stochastic ascent methods is driven by the objective function and how the ascent directions are estimated.
-First, the objective function shall be (pseudo) concave to find its global maximum (Bottou, 1998).
-Second, the convergence rate is influenced by the distribution of the stochastic ascent estimates (Chen & Luss, 2018; Ajalloeian & Stich, 2020).
+First, the objective function shall be (pseudo) concave to find its global maximum (Bottou,, 1998).
+Second, the convergence rate is influenced by the distribution of the stochastic ascent estimates (Chen and Luss,, 2018; Ajalloeian and Stich,, 2020).
 In this paper, we rigorously study policy-gradient methods with exploration-based reward shaping through the lens of these two optimization theory aspects.
@@
 In Section 2, we provide the background about policy gradients and exploration.
-Section 3 focuses on the effect of exploration on the learning objective while Section 4 is dedicated to the effect on the gradient estimates used in the policy-gradient algorithms [^1] Experimental details and implementations can be found at
-https://github.com/adrienBolland/micro-rl-lib.
-. Conclusions and future works are discussed in Section 5.
+Section 3 focuses on the effect of exploration on the learning objective while Section 4 is dedicated to the effect on the gradient estimates used in the algorithms.
+Both sections include illustrations on pathological examples [^1] Implementations can be found at https://github.com/adrienBolland/micro-rl-lib..
+Conclusions and future works are discussed in Section 5.
 2 Background
-In this section, we introduce Markov decision processes and policy gradients with intrinsic exploration.
+Let us introduce Markov decision processes and policy gradients with intrinsic exploration.
 2.1 Markov Decision Processes
-We study problems in which an agent makes sequential decisions in a stochastic environment (Sutton & Barto, 2018).
-The environment is modeled with an infinite-time Markov decision process (MDP) composed of a state space $\mathcal{S}$, an action space $\mathcal{A}$, an initial state distribution with density $p_{0}$, a transition distribution (modeling the dynamics) with conditional density $p$, a bounded reward function $\rho$, and a discount factor $\gamma\in[0,1)$.
-When an agent interacts with the MDP, first, an initial state $s_{0}\sim p_{0}(\cdot)$ is sampled, then, the agent provides at each time step $t$ an action $a_{t}\in\mathcal{A}$ leading to a new state $s_{t+1}\sim p(\cdot|s_{t},a_{t})$.
-Such a sequence of states and actions $h_{t}=(s_{0},a_{0},\dots,s_{t-1},a_{t-1},s_{t})\in\mathcal{H}$ is called a history and $\mathcal{H}$ is the set of all histories of any arbitrary length.
-In addition, after an action $a_{t}$ is executed, a reward $r_{t}=\rho(s_{t},a_{t})\in\mathbb{R}$ is observed.
-A policy $\pi\in\Pi=\mathcal{S}\rightarrow\mathcal{P}(\mathcal{A})$ is a mapping from the state space $\mathcal{S}$ to the set of probability measures on the action space $\mathcal{P}(\mathcal{A})$, where $\pi(a|s)$ is the associated conditional probability density of action $a$ in state $s$.
-The function $J:\Pi\rightarrow\mathbb{R}$ is defined as the function mapping any policy $\pi$ to the expected discounted sum of rewards gathered by an agent interacting in the MDP by sampling actions from the policy $\pi$.
+We study problems in which an agent makes sequential decisions in a stochastic environment (Sutton and Barto,, 2018).
+The environment is modeled as an infinite-time Markov decision process (MDP) composed of a state space $\mathcal{S}$, an action space $\mathcal{A}$, an initial state distribution with density $p_{0}$, a transition distribution (modeling the dynamics) with conditional density $p$, a bounded reward function $\rho$, and a discount factor $\gamma\in[0,1)$.
+When an agent interacts with the MDP, first an initial state $s_{0}\sim p_{0}(\cdot)$ is sampled, then, at each time step $t$, the agent provides an action $a_{t}\in\mathcal{A}$, leading to a new state $s_{t+1}\sim p(\cdot|s_{t},a_{t})$.
+Finally, a reward $r_{t}=\rho(s_{t},a_{t})\in\mathbb{R}$ is observed after each action $a_{t}$.
+A policy $\pi\in\Pi=\mathcal{S}\rightarrow\mathcal{P}(\mathcal{A})$ is a mapping from the state space $\mathcal{S}$ to the set of probability measures on the action space $\mathcal{P}(\mathcal{A})$, where $\pi(a|s)$ is the associated conditional probability of action $a$ in state $s$.
+The function $J:\Pi\rightarrow\mathbb{R}$ is defined as a function mapping any policy $\pi$ to the expected discounted sum of rewards gathered by an agent interacting with the MDP by sampling actions from $\pi$.
 We call return of the policy $\pi$ the value provided by that function
 $J(\pi)=\frac{1}{1-\gamma}\underset{\begin{subarray}{c}s\sim d^{\pi,\gamma}(\cdot)\\ a\sim\pi(\cdot|s)\end{subarray}}{\mathbb{E}}\left[\rho(s,a)\right]\;,$ (1)
-where $d^{\pi,\gamma}(\cdot)$ is the discounted state-visitation probability (Manne, 1960).
-In reinforcement learning, we seek to find an optimal policy $\pi^{*}$ maximizing the expected discounted sum of rewards $J$.
+where $d^{\pi,\gamma}(s)=(1-\gamma)\sum_{t=0}^{\infty}\gamma^{t}p_{t}^{\pi}(s)$ is the state-visitation measure $\forall s\in\mathcal{S}$, and $p_{t}^{\pi}$ is the state probability after $t$ time steps following the policy $\pi$.
+In reinforcement learning, an optimal policy $\pi^{*}\in\Pi$ maximizes the expected discounted sum of rewards $J$.
 2.2 Policy-Gradient Algorithms
-Policy-gradient algorithms (locally) optimize a parameterized policy $\pi_{\theta}$ to find the optimal parameter $\theta^{*}$ for which the return of the policy $J(\pi_{\theta^{*}})$ is maximized.
-Naively maximizing the return may provide sub-optimal results.
-This problem is mitigated in practice with exploration strategies, which consist in optimizing a surrogate learning objective $L$ that intrinsically encourages certain behaviors.
-In this work, we consider reward-shaping strategies where the expected discounted sum of rewards is extended by $K$ additional reward terms $\rho_{i}^{int}$, called intrinsic motivation terms, and optimize the learning objective
-$L(\theta)$ $=\frac{1}{1-\gamma}\underset{\begin{subarray}{c}s\sim d^{\pi_{\theta},\gamma}(\cdot)\\ a\sim\pi_{\theta}(\cdot|s)\end{subarray}}{\mathbb{E}}\left[\rho(s,a)+\sum_{i=0}^{K-1}\lambda_{i}\rho_{i}^{int}(s,a)\right]$
-$=J(\pi_{\theta})+J^{int}(\pi_{\theta})\;,$ (2)
-where $\lambda_{i}$ are non-negative weights for each intrinsic reward and where $J^{int}(\pi_{\theta})$ is the intrinsic return of the policy.
+Policy-gradient algorithms (locally) optimize a parameterized policy $\pi_{\theta}$ to find the parameter $\theta^{*}$ that maximizes the return of the policy.
+Naively maximizing the return may provide suboptimal results.
+This problem is mitigated in practice by exploration strategies, which typically consist of optimizing surrogate learning objectives that encourage certain behaviors.
+In this work, we consider reward-shaping strategies in which the expected discounted sum of rewards is extended by $K$ additional reward terms $\rho_{k}^{int}$, called intrinsic motivation terms, and we optimize the learning objective
+$L(\theta)$ $=\frac{1}{1-\gamma}\underset{\begin{subarray}{c}s\sim d^{\pi_{\theta},\gamma}(\cdot)\\ a\sim\pi_{\theta}(\cdot|s)\end{subarray}}{\mathbb{E}}\left[\rho(s,a)+\sum_{k=0}^{K-1}\lambda_{k}\rho_{k}^{int}(s,a)\right]=J(\pi_{\theta})+J^{int}(\pi_{\theta})\;,$ (2)
+where $\lambda_{k}$ are non-negative weights for each intrinsic reward, and $J^{int}(\pi_{\theta})$ is the intrinsic return of the policy.
 The parameter maximizing the learning objective is denoted by $\theta^{\dagger}$, which we distinguish from the optimal policy parameter $\theta^{*}$.
-Most of the intrinsic motivation terms can be classified in the two following groups.
+Most of the intrinsic motivation terms can be classified into the following two groups.
 Uncertainty-based motivations.
-It is common to provide bonuses for performing actions that reduce the uncertainty of the agent about its environment (Pathak et al., 2017; Burda et al., 2018; Zhang et al., 2021c).
-The intrinsic motivation terms are then proportional to the prediction errors of a model of the MDP dynamics.
-The latter model is usually learned.
+It is common to provide bonuses for performing actions that reduce the agent uncertainty about an internal environment model (Pathak et al.,, 2017; Burda et al.,, 2018; Zhang et al., 2021c, ).
+The intrinsic motivation terms are then proportional to the prediction errors of that model, which is typically learned.
 Entropy-based motivations.
-It is also common to provide bonuses for visiting states and/or playing actions that are less likely in histories (Haarnoja et al., 2019; Hazan et al., 2019).
-In this work, we focus on two of these bonuses
-$\rho^{s}(s,a)$ $=-\log d^{\pi_{\theta},\gamma}(\phi(s))$ (3)
-$\rho^{a}(s,a)$ $=-\log\pi_{\theta}(a|s)\;,$ (4)
-where $\phi(s)$ is a feature built from the state $s$.
-The corresponding intrinsic returns are maximized for policies that visit uniformly every feature, and for policies with uniformly distributed actions in each state, respectively.
-Note that these rewards require to estimate the distribution over the states and/or actions.
-Furthermore, they implicitly depend on the policy parameter $\theta$.
-The second technique is usually referred to as entropy regularization.
-In this work, we consider on-policy policy-gradient algorithms, which were among others reviewed by (Duan et al., 2016) and (Andrychowicz et al., 2020).
+It is also common to motivate agents to visit states and/or select actions that are less likely.
+It is typically achieved using intrinsic rewards that depend on the policy (Williams and Peng,, 1991; Haarnoja et al.,, 2019) or on the state-visitation measure (Bellemare et al.,, 2016; Hazan et al.,, 2019; Lee et al.,, 2019; Islam et al.,, 2019; Guo et al.,, 2021; Zhang et al., 2021b, )
+$\rho^{a}(s,a)$ $=-\log\pi_{\theta}(a|s)$ $J^{a}(\pi_{\theta})\propto\mathbb{E}_{s\sim d^{\pi_{\theta},\gamma}(\cdot)}\left[\mathcal{H}_{a}\left(\pi_{\theta}(a|s)\right)\right]$ (3)
+$\rho^{s}(s,a)$ $=-\log d^{\pi_{\theta},\gamma}(z)$ $J^{s}(\pi_{\theta})\propto\mathcal{H}_{z}\left(d^{\pi_{\theta},\gamma}(z)\right)\;,$ (4)
+where $z=\phi(s)$ is a feature built from the state $s$, and where $\mathcal{H}_{x}(q(x))$ is the entropy of $q$.
+The corresponding intrinsic returns, $J^{a}(\pi_{\theta})$ and $J^{s}(\pi_{\theta})$, are maximized for policies with actions uniformly distributed in each state, and for policies that visit every feature uniformly, respectively.
+Importantly, these rewards require estimating the distribution over the states and/or actions, which implicitly depends on $\theta$.
+Notice that the first technique is usually referred to as entropy regularization.
+In this work, we consider on-policy policy-gradient algorithms, which were, among others, reviewed by Duan et al., (2016) and Andrychowicz et al., (2020).
 These algorithms optimize differentiable parameterized policies with gradient-based local optimization.
-They iteratively approximate an ascent direction $\hat{d}$ relying on histories sampled from the policy in the MDP and update the parameters in the ascent direction, or in a combination of the previous ascent directions (Hinton et al., 2012; Kingma & Ba, 2014).
-For the sake of simplicity and without loss of generality, we consider that the ascent direction $\hat{d}$ is composed of the sum of an estimate of the gradient of the return $\hat{g}\approx\nabla_{\theta}J(\pi_{\theta})$ and an estimate of the gradient of the intrinsic return $\hat{i}\approx\nabla_{\theta}J^{int}(\pi_{\theta})$.
-In practice, the first is usually unbiased while the second is computed neglecting some partial derivatives of $\theta$ and is thus biased, typically neglecting the influence of the policy on the intrinsic reward.
+They iteratively approximate an ascent direction $\hat{d}$, relying on samples from the policy in the MDP, and update the parameters in that ascent direction, or in a combination of previous ascent directions (Hinton et al.,, 2012; Kingma and Ba,, 2014).
+For the sake of simplicity and without loss of generality, we consider that the ascent direction $\hat{d}$ is the sum of an estimate of the gradient of the return $\hat{g}\approx\nabla_{\theta}J(\pi_{\theta})$ and an estimate of the gradient of the intrinsic return $\hat{i}\approx\nabla_{\theta}J^{int}(\pi_{\theta})$.
+In practice, the first may be unbiased, while the second is computed by neglecting some partial derivatives of $\theta$ and is thus biased, typically neglecting the influence of the policy on the intrinsic reward.
 3 Study of the Learning Objective
 In this section, we study the influence of the exploration terms on the learning objective defined in equation (2).
-We define two criteria under which the learning objective can be globally optimized by ascent methods, and such that the solution is close to an optimal policy.
+We define two criteria under which the learning objective can be globally optimized by ascent methods and under which the solution is close to an optimal policy.
 We then graphically illustrate how exploration modifies the learning objective to remove local extrema.
 3.1 Policy-Gradient Learning Objective
-Policy-gradient algorithms using exploration maximize the learning objective function $L$, as defined in equation (2).
-We introduce two criteria related to this learning objective for studying the performance of the policy-gradient algorithm.
+Policy-gradient algorithms using exploration maximize the learning objective function $L$ as defined in equation (2).
+We introduce two criteria related to this learning objective to study the performance of the policy-gradient algorithm.
 First, we say that a learning objective $L$ is $\epsilon$-coherent when its global maximum is in an $\epsilon$-neighborhood of the return of an optimal policy.
@@
 $J(\pi_{\theta^{*}})-J(\pi_{\theta^{\dagger}})\leq\epsilon\;,$ (5)
-where $\theta^{*}\in\text{argmax}_{\theta}J(\pi_{\theta})$ and where $\theta^{\dagger}\in\text{argmax}_{\theta}L(\theta)$.
+where $\theta^{*}\in\arg\max_{\theta}J(\pi_{\theta})$ and where $\theta^{\dagger}\in\arg\max_{\theta}L(\theta)$.
 Pseudoconcavity criterion. A learning objective $L$ is pseudoconcave if, and only if,
 $\exists!\>\theta^{\dagger}:\nabla L(\theta^{\dagger})=0\land L(\theta^{\dagger})=\max_{\theta}L(\theta)\;.$ (6)
-If the pseudoconcavity criterion is respected, there is a single optimum, and it is thus possible to globally optimize the learning objective function by (stochastic) gradient ascent (Bottou, 2010) [^2] For the sake of keeping discussions simple, the definition of pseudoconcavity is simplified (Mangasarian, 1975), and additional assumptions on the stochastic gradient estimates are neglected..
+If the pseudoconcavity criterion is satisfied, there is a single optimum, and it is thus possible to globally optimize the learning objective function by (stochastic) gradient ascent (Bottou,, 2010) [^2] The literature on policy-gradient algorithms usually relies on the Polyak–Łojasiewicz inequality and gradient dominance (Polyak,, 1963; Lojasiewicz,, 1963; Karimi et al.,, 2016) rather than pseudoconvexity (Mangasarian,, 1975), all of which imply that every stationary point is a global minimum (invexity), from which global convergence of gradient descent methods follows (Ben-Israel and Mond,, 1986; Mishra and Giorgi,, 2008; Bottou,, 2010).
+Gradient dominance furthermore ensures linear rates of convergence (Karimi et al.,, 2016).
+Under quasiconvexity, invexity and pseudoconvexity coincide (Mishra and Giorgi,, 2008) and are thus implied by gradient dominance.
+For the sake of keeping the discussion simple, the definitions of pseudoconvexity and pseudoconcavity are simplified, and additional assumptions on the stochastic gradient estimates are neglected..
 If the learning objective is furthermore $\epsilon$-coherent, the latter solution is also a near-optimal policy, where $\epsilon$ is the bound on the suboptimality of its return.
-Let us finally remind a theorem from Ng et al. (1999).
+Let us finally recall a theorem of Ng et al., (1999).
 Consistency Theorem.
-The learning objective $L$ is $\epsilon$-coherent, with $\epsilon=0$, in any MDP with state space $\mathcal{S}$, action space $\mathcal{A}$ and factor $\gamma$, if, and only if, $J(\theta)=L(\theta)$ for all $\theta$.
-The intrinsic rewards are furthermore potential based.
-This theorem states that there is no MDP-agnostic exploration method that guarantees consistency with $\epsilon$ equal to zero and that modifies the objective function.
-This type of exploration is only possible with potential-based reward shaping (Ng et al., 1999).
-In conclusion, if the return is not pseudoconcave, there is a trade-off between the two criteria, which can not be resolved by potential-based exploration.
+The learning objective $L$ is $\epsilon$-coherent, with $\epsilon=0$, in any MDP with state space $\mathcal{S}$, action space $\mathcal{A}$ and discount factor $\gamma$, if, and only if, $J(\pi_{\theta})=L(\theta)$ for all $\theta$.
+Furthermore, the intrinsic rewards are potential based.
+This theorem establishes a trade-off between the coherence and pseudoconcavity criteria.
+Either an exploration method guarantees coherence with $\epsilon=0$ in any MDP, but then the learning objective equals the return and the pseudoconcavity criterion is violated when the return is not pseudoconcave itself.
+Otherwise, the pseudoconcavity criterion is satisfied, but then the exploration method is MDP-dependent or coherence is guaranteed at best with $\epsilon>0$.
 3.2 Illustration of the Effect of Exploration on the Learning Objective
-Exploration is of paramount importance in environments with complex dynamics and reward functions, where many locally optimal policies may exist (Lee et al., 2019; Liu & Abbeel, 2021; Zhang et al., 2021b).
+Exploration is of paramount importance in environments with complex dynamics and reward functions, where many locally optimal policies may exist (Lee et al.,, 2019; Liu and Abbeel,, 2021; Zhang et al., 2021b, ).
 In the following, we first define such an environment and a policy parameterization that will serve as an example to illustrate the effect of exploration on the optimization process.
-For the sake of the analysis, we then represent the learning objectives associated to different exploration strategies, and depict their global and local optima.
-Learning objectives with a single optimum respect the pseudoconcavity criterion.
-In addition, we represent the neighborhood $\Omega$ of the optimal policy parameters, such that any learning objective with its global maximum within this region is coherent for a given $\epsilon$.
-In light of the coherence and the pseudoconcavity criteria, we finally elaborate on the policy parameter computed by stochastic gradient ascent algorithms.
-We consider the environment illustrated in Figure 1(a) where a car moves in a valley (Bolland et al., 2023).
-We denote by $x$ and $v$ the position and speed of the car, both composing its state $s=(x,v)$.
-The valley contains two separate low points, positioned in $x_{initial}=-3$ and $x_{target}=3$, separated by a peak.
-The car starts at rest $v_{0}=0$ at the highest low point $x_{0}=x_{initial}$ and receives rewards proportional to the depth of the valley at its current position.
-The reward function is provided in Figure 1(b).
-We consider a policy $\pi_{K,\sigma}(a|s)=\mathcal{N}(a|\mu_{K}(s),\sigma)$, namely a normally disturbed proportional controller with $\mu_{K}(s)=K\times(x-x_{target})$, parameterized by the vector $\theta=(K,\sigma)$.
-Figure 1(c) illustrates the contour map of the return of the policy as a function of the parameters $K$ and $\sigma$.
-The optimal parameters are represented by black dots and correspond to policies that drive the car to pass the peak and reach the lowest valley point in $x_{target}$.
-The green area represents the set of parameters $\Omega=\left\{\theta^{\prime}|\max_{\theta}J(\pi_{\theta})-J(\pi_{\theta^{\prime}})\leq\epsilon\right\}$ for $\epsilon=1$.
+For the sake of the analysis, we then represent the learning objectives associated with different exploration strategies, and depict their global and local optima.
+Learning objectives with a single optimum satisfy the pseudoconcavity criterion.
+In addition, we represent the neighborhood $\Omega$ of the optimal policy parameters such that any learning objective with its global maximum within this region is coherent for a given $\epsilon$.
+In light of the coherence and the pseudoconcavity criteria, we finally elaborate on the policy parameters computed by stochastic gradient ascent algorithms.
+We consider the environment illustrated in Figure 1(a), where a car moves in a valley (Bolland et al.,, 2023).
+We denote by $x$ and $v$ the position and speed of the car, which together compose its state $s=(x,v)$.
+The valley contains two valley floors, located at $x_{initial}=-3$ and $x_{target}=3$, separated by a peak.
+The car starts at rest $v_{0}=0$ at the higher floor $x_{0}=x_{initial}$ and receives rewards proportional to the depth of the valley at its current position.
+The reward function is shown in Figure 1(b).
+We consider a policy $\pi_{K,\sigma}(a|s)=\mathcal{N}(a|\mu_{K}(s),\sigma)$ with $\mu_{K}(s)=K\times(x-x_{target})$, parameterized by the vector $\theta=(K,\sigma)$.
+It corresponds to a noisy proportional controller.
+Figure 1(c) illustrates the contour map of the return, annotated with local optima, and shows in green the set of near-optimal policies.
+The optimal policy drives the car to reach the lowest floor in $x_{target}$.
 (a) Hill environment.
@@
 In Figure 1(c), the return of the policy $\pi_{K,\sigma}$ with the global and local maximum represented in black and grey, together with their respective return values.
+In green $\Omega=\left\{\theta^{\prime}|\max_{\theta}J(\pi_{\theta})-J(\pi_{\theta^{\prime}})\leq\epsilon=1\right\}$.
 Figure 2 illustrates the learning objective with the intrinsic rewards $\rho^{s}(s,a)=-\log d^{\pi_{K,\sigma},\gamma}(\phi(s))$, from equation (4), and $\rho^{a}(s,a)=-\log\pi_{K,\sigma}(a|s)$, from equation (3), for different values of the corresponding weights $\lambda_{s}$ and $\lambda_{a}$.
 Here, the feature is the position in the valley $\phi(s)=x$.
-First, we observe that for weights approaching zero, the parameter $\theta^{\dagger}$ maximizing the learning objective, represented by a black dot, corresponds to a policy with a high return.
-More precisely, it is in the green set $\Omega$ such that $\epsilon$-coherence is guaranteed for the value $\epsilon=1$.
-Larger weights require larger values of $\epsilon$ for guaranteeing the $\epsilon$-coherence criterion.
-Nevertheless, when increasing the weights, we observe that the learning objective eventually becomes pseudoconcave.
-There appears to be a trade-off between the two criteria.
-In Figure 2(b), we observe that in this environment, there is a learning objective that respects the pseudoconcavity criterion and the $\epsilon$-coherence criterion for $\epsilon=1$.
-Indeed, there is a single global maximum in Figure 2(b) represented by a black dot that is furthermore part of the set $\Omega$.
-Shaping the reward function with an exploration strategy based on the state-visitation entropy appears to be a good solution for optimizing the policy.
-However, a notable drawback is that the reward depends on the policy and its (gradient) computation requires to estimate a complex probability measure.
-In this example, the intrinsic reward function itself was estimated by Monte-Carlo sampling for every parameter, which would not scale for complex problems and requires approximations and costly evaluation strategies (Islam et al., 2019).
-In Appendix A we present an alternative problem-dependent intrinsic reward, independent of the policy parameters and thus simple to compute efficiently, that still respects the pseudoconcavity and $\epsilon$-coherence criteria, and in Appendix B we extend the study to more complex environments from the MiniGrid library (Chevalier-Boisvert et al., 2023) where the policy is a deep neural network and the state-visitation probability is approximated.
+On the one hand, the smaller the exploration weights, the closer the parameter $\theta^{\dagger}$ maximizing the learning objective is to $\theta^{*}$, and the smaller $\epsilon$ for which coherence is guaranteed.
+Furthermore, the return of the policy corresponding to $\theta^{\dagger}$ is then also higher.
+In the particular case depicted in the figure, we observe that, only for the smallest weights, the parameter $\theta^{\dagger}$ maximizing the learning objective, represented by a black dot, is inside the green set $\Omega$, such that $\epsilon$-coherence is guaranteed for $\epsilon=1$.
+On the other hand, as the weights increase, we observe that the learning objective eventually becomes pseudoconcave.
+As explained, this comes at the cost of a higher $\epsilon$ for guaranteeing $\epsilon$-coherence; there is thus a trade-off between the two criteria.
+In Figure 2(b), we observe that in this environment, there is a learning objective that satisfies both the pseudoconcavity criterion and the $\epsilon$-coherence criterion for $\epsilon=1$.
+Indeed, there is a single global maximum represented by a black dot, which is part of $\Omega$.
 (a) $\lambda_{s}=0.05$ and $\lambda_{a}=0$
@@
 The darker the map, the larger the learning objective value.
-The green area represents the set $\Omega=\left\{\theta^{\prime}|\max_{\theta}J(\pi_{\theta})-J(\pi_{\theta^{\prime}})\leq\epsilon=1\right\}$, such that when the parameter maximizing the learning objective is part of $\Omega$, then the learning objective function is $\epsilon$-coherent with $\epsilon=1$.
 The black dot is the parameter $\theta^{\dagger}$ globally maximizing the learning objective and the grey dot is the local (non-global) maximum of the learning objective if it exists.
 Both are labeled with the return values of the corresponding policies.
-The observations suggest that well-chosen exploration strategies can lead to learning objective functions that satisfy the two criteria defined in the previous section, thereby guaranteeing that policies suboptimal by at most $\epsilon$ can be computed by local optimization.
-When designing exploration strategies, it is essential to keep in mind that we modify the learning objective for the algorithms to converge to optimal policy parameters, which can be achieved when both criteria are respected.
-While strategies such as enforcing entropy can be effective in some environments, they are only heuristic strategies and not to be relied upon exclusively.
+The green area represents the set $\Omega=\left\{\theta^{\prime}|\max_{\theta}J(\pi_{\theta})-J(\pi_{\theta^{\prime}})\leq\epsilon=1\right\}$.
+Shaping the reward function with an exploration strategy based on the state-visitation entropy appears to be a good solution for optimizing the policy.
+However, a notable drawback is that the reward depends on the policy, and its (gradient) computation requires estimating a complex probability measure.
+In this example, the intrinsic reward function itself was estimated by Monte Carlo sampling for each parameter, which would not scale for complex problems and requires approximations and costly evaluation strategies.
+In Appendix A, we extend the study to more complex environments, where the policy is a neural network and the state-visitation probability is approximated.
+The observations suggest that well-chosen exploration strategies can lead to learning objective functions that satisfy the two criteria defined in the previous section, thereby guaranteeing that policies that are suboptimal by at most $\epsilon$ can be computed by local optimization.
+When designing exploration strategies, it is essential to keep in mind that we modify the learning objective so that the algorithms can converge to optimal policy parameters, which can be achieved when both criteria are satisfied.
+While strategies such as enforcing entropy can be effective in some environments, they are only heuristic strategies and should not be relied upon exclusively.
 Furthermore, as illustrated, both criteria may be subject to a trade-off.
-In more complex environments, an efficient exploration strategy may require to balance both criteria, e.g., through a schedule on the learning objective weights.
+In more complex environments, an efficient exploration strategy may require balancing both criteria, e.g., through a schedule on the learning objective weights.
 4 Study of the Ascent Direction Distribution
-Optimizing pseudoconcave functions with stochastic ascent methods are guaranteed to converge (at a certain rate) under assumptions on the distribution of the gradient estimates at hand (Bottou, 2010; Chen & Luss, 2018; Ajalloeian & Stich, 2020).
+Optimizing pseudoconcave functions with stochastic ascent methods is guaranteed to converge (at a certain rate) under assumptions about the distribution of the gradient estimates (Bottou,, 2010; Chen and Luss,, 2018; Ajalloeian and Stich,, 2020).
 In this section, we study the influence of the exploration terms on this distribution in the context of policy gradients.
-More precisely, we study the probability of improving the learning objective and the return with stochastic ascent steps.
-Intuitively, they shall be sufficiently large for the algorithm to be efficient.
-We formalize this intuition and illustrate how exploration strategies can increase these probabilities, leading to more efficient algorithms.
+More precisely, we study the probabilities of improving the learning objective and the return through stochastic ascent steps.
+Intuitively, these probabilities should be sufficiently large for the algorithm to be efficient.
+We formalize this intuition and illustrate how exploration can increase them, leading to more efficient algorithms.
 4.1 Policy-Gradient Estimated Ascent Direction
-In general, gradient ascent algorithms update parameters in a direction $\hat{d}$ in order to locally improve an objective function $f$.
-The quality of these algorithms can therefore be studied (for a small step size $\alpha\rightarrow 0$) through the random variable representing the quantity by which the objective increases for each $\theta$
-$X$ $=f(\theta+\alpha\hat{d})-f(\theta)=\alpha\>\langle\hat{d},\nabla_{\theta}f(\theta)\rangle\>,$ (7)
-[…]
-The (asymptotic) convergence of a gradient ascent algorithm is usually studied by bounding the expectation of $X$.
-Such bounds depend, among others, on the expectation of $\hat{d}$, which equals $\nabla_{\theta}f(\theta)$ when unbiased, and depend on the variance of $\hat{d}$, which deteriorates the expected convergence rate.
-Finding rates depending on the algorithm is an active field of research.
-In parallel, a slightly more general problem is to quantify if the expectation of $X$ is driven by rare events.
-Intuitively, an algorithm with unbiased gradient estimates has positive expected improvements $\mathbb{E}[X]>0$, and should theoretically converge, but may be inefficient in practice if positive events $X>0$ rarely occur.
-We illustrate in the next section that this phenomenon makes reinforcement learning in sparse-reward environments particularly hard.
-To the best of our knowledge, no existing result fits to the study of policy gradients.
-We therefore introduce two new criteria on the probability of improvement $P(X>0)$, which we empirically validate afterwards.
-First, we define an exploration strategy as $\delta$-efficient if, and only if, following the ascent direction $\hat{d}\approx\nabla_{\theta}L(\theta)$ has a probability at least $\delta$ to increasing the learning objective $L(\theta)$ almost everywhere.
-Second, an exploration strategy is $\delta$-attractive if, and only if, there exists a neighborhood of $\theta^{\dagger}$ containing the parameter $\theta^{int}$ maximizing the intrinsic return $J^{int}$, where the probability of increasing the return by following $\hat{d}$ is almost everywhere at least equal to $\delta$.
-Note that each probability measure and random variable is a function of $\theta$, which we do not explicitly write for the sake of keeping notations simple.
+In general, ascent algorithms update parameters in a direction $\hat{d}$ to locally improve an objective function $f$.
+The quality of these algorithms can therefore be studied (for a small step size $\alpha\rightarrow 0$) through the random variable representing the amount by which the objective increases for each $\theta$
+$X$ $=f(\theta+\alpha\hat{d})-f(\theta)=\alpha\>\langle\hat{d},\nabla_{\theta}f(\theta)\rangle+O(\alpha^{2})\>,$ (7)
+where $\langle\cdot,\cdot\rangle$ is the Euclidean scalar product. This variable depends on the random event $\hat{d}$.
+The convergence of ascent algorithms is usually studied by bounding the suboptimality of $f(\theta)$ as a function of the number of ascent steps.
+In this active field of research, the rates are characterized by assumptions on $\hat{d}$ or, equivalently, on $X$.
+Depending on the assumptions, the efficiency of algorithms is, for example, studied as a function of the statistical moments of the ascent estimate or of the probability of alignment between the gradient and the estimate.
+We choose to work in the context in which gradient estimates have positive expected improvements $\mathbb{E}[X]>0$ (e.g., because the estimates are unbiased), and we quantify whether this expectation is positive due to rare realizations of large values of $X$.
+Intuitively, an algorithm with positive expected improvements $\mathbb{E}[X]>0$ should theoretically converge but may be inefficient in practice if positive events $X>0$ rarely occur.
+Below, we study the influence of such events, typically observed in sparse-reward environments and for which the chosen context is sufficient to assess efficiency without more assumptions.
+Further work may proceed differently to derive convergence rates or quantify other phenomena.
+In the following, we introduce two new criteria on the probability of improvement $P(X>0)$, which we empirically validate later.
+First, we define an exploration strategy as $\delta$-efficient if, and only if, following the ascent direction $\hat{d}\approx\nabla_{\theta}L(\theta)$ has a probability of at least $\delta$ of increasing the learning objective $L(\theta)$ almost everywhere (within a region of the parameter space).
+Second, an exploration strategy is $\delta$-attractive if, and only if, there exists a neighborhood of $\theta^{\dagger}$ containing the parameter $\theta^{int}$ that maximizes the intrinsic return $J^{int}$, where the probability of increasing the return by following $\hat{d}$ is almost everywhere at least equal to $\delta$.
+Note that each probability measure and random variable are functions of $\theta$, which we do not explicitly write to keep the notation simple.
 Efficiency criterion. An exploration strategy is $\delta$-efficient if, and only if,
@@
 where $\theta^{int}=\text{argmax}_{\theta}J^{int}(\pi_{\theta})$, $B(\theta^{\dagger})$ is a ball centered in $\theta^{\dagger}$, and $G=\>\langle\hat{d},\nabla_{\theta}J(\pi_{\theta})\rangle$.
-The efficiency criterion quantifies how often a stochastic gradient ascent step improves the learning objective.
+The efficiency criterion quantifies how often a stochastic gradient ascent step improves the learning objective (in a region of the parameter space).
 The larger, the better the learning objective and its stochastic ascent direction approximations.
-The rationale behind the attraction criterion is that in many exploration strategies, the intrinsic reward is dense, and it is then presumably easy to optimize the intrinsic return in the sense that $\mathbb{P}(\>\langle\hat{i},\nabla_{\theta}J^{int}(\pi_{\theta})\rangle>0)$ is large.
+The rationale behind the attraction criterion is that, in many exploration strategies, the intrinsic reward is dense, and it is then presumably easy to optimize the intrinsic return in the sense that $\mathbb{P}(\>\langle\hat{i},\nabla_{\theta}J^{int}(\pi_{\theta})\rangle>0)$ is large.
 It implies that it is easy to locally improve the learning objective by (solely) increasing the value of the intrinsic motivation terms.
-It furthermore implies that policy-gradient algorithms may be subject to converging towards $\theta^{int}$ rather than $\theta^{\dagger}$ when $\mathbb{P}(\>\langle\hat{d},\nabla_{\theta}J(\pi_{\theta})\rangle>0)$ is small.
-If the criterion is respected for large $\delta$, the latter is less likely to happen as policy gradients will eventually tend to improve the return of the policy if the parameter approaches $\theta^{int}$ and enters the ball $B(\theta^{\dagger})$; eventually converging towards $\theta^{\dagger}$.
+It furthermore implies that policy-gradient algorithms may tend to converge towards $\theta^{int}$ rather than $\theta^{\dagger}$ when $\mathbb{P}(\>\langle\hat{d},\nabla_{\theta}J(\pi_{\theta})\rangle>0)$ is small.
+If the attraction criterion is satisfied for large $\delta$, the latter is less likely to happen as policy gradients will eventually tend to improve the return of the policy if the parameter approaches $\theta^{int}$ and enters the ball $B(\theta^{\dagger})$; eventually converging towards $\theta^{\dagger}$.
 These two new criteria on $\hat{d}$ are independent of the previous ones on $L$, which only captured the quality of the deterministic learning objective functions.
-In the particular cases where the learning objectives $L$ are $\epsilon$-coherent, for $\epsilon=0$, and pseudoconcave, e.g., with potential-based intrinsic rewards, only the distribution of estimates $\hat{d}$ can explain why some algorithms succeed and others fail.
-Finally, the value of $\delta$ in the new criteria we introduce can be related to the variance of the estimate $\hat{d}$ under some assumptions, e.g., with Cantelli’s concentration inequalities.
+In the particular cases where the learning objectives $L$ are $\epsilon$-coherent, for $\epsilon=0$, and pseudoconcave, e.g., with potential-based intrinsic rewards, only the distribution of estimates $\hat{d}$ can explain why some algorithms outperform others.
+Finally, the value of $\delta$ in the two new criteria could be related to the variance of the estimate $\hat{d}$, for example via Cantelli’s concentration inequalities.
 4.2 Illustration of the Effect of Exploration on the Estimated Ascent Direction
-Exploration is usually promoted and tested for problems where the reward function is sparse, typically in maze-environments (Islam et al., 2019; Liu & Abbeel, 2021; Guo et al., 2021).
-In this section, we first introduce a new maze-environment with sparse rewards where we illustrate the influence of exploration on the gradient estimates of the learning objective.
-To this end, we present two learning objective functions and elaborate on the influence of exploration on the performance of policy-gradient algorithms in the light of the efficiency and attraction criteria.
-Let us consider a maze-environment consisting of a horizontal corridor composed of $S\in\mathbb{N}$ tiles.
-The state of the environment is the index of the tile $s\in\left\{1,\dots,S\right\}$, and the actions consist in going left $a=-1$ or right $a=+1$.
-When an action is taken, the agent stays idle with probability $p=0.7$, and moves with probability $0.31-p=$ 0.3 in the direction indicated by the action, then $s^{\prime}=\min(S,\max(1,s+a))$.
-The agent starts in state $s=1$ and the target state $s=S=15$ is absorbing.
-Zero rewards are observed except when the agent reaches the target state where a reward $r=100$ is observed.
+Exploration is usually promoted and tested for problems where the reward function is sparse, typically in maze environments (Islam et al.,, 2019; Liu and Abbeel,, 2021; Guo et al.,, 2021).
+In this section, we first introduce a new maze environment with sparse rewards to illustrate the influence of exploration on the gradient estimates of the learning objective.
+To this end, we present two learning objective functions and elaborate on the influence of exploration on the performance of policy-gradient algorithms in light of the efficiency and attraction criteria.
+Let us consider a maze environment consisting of a horizontal corridor composed of $S\in\mathbb{N}$ tiles.
+The state of the environment is the index of the tile $s\in\left\{1,\dots,S\right\}$, and the actions consist of going left a=1a=-1 or right $a=+1$.
+When an action is taken, the agent stays idle with probability $p=0.7$, and moves with probability 1p=0.31-p=0.3 in the direction indicated by the action, then $s^{\prime}=\min(S,\max(1,s+a))$.
+The agent starts in state $s=1$, and the target state $s=S=15$ is absorbing.
+A zero reward is observed except when the agent reaches the target state, where a reward of $r=100$ is observed.
 A discount factor of $\gamma=0.99$ is considered.
-Finally, we study the policy going with probability $\theta$ to the right and probability $1-\theta$ to the left, and with density
-$\pi_{\theta}(a|s)=\left\{\begin{array}[]{cl}\theta&\quad\textrm{if }a=1\\ 1-\theta&\quad\textrm{if }a=-1\;.\end{array}\right.$ (13)
+Finally, we study the policy that moves right with probability $\theta$ and left with probability $1-\theta$, with density
+$\pi_{\theta}(a|s)=\left\{\begin{array}[]{cl}\theta&\quad\textrm{if }a=+1\\ 1-\theta&\quad\textrm{if }a=-1\;.\end{array}\right$. (13)
 The return $J(\pi_{\theta})$ is represented in black in Figure 3(a) as a function of $\theta$ along with two intrinsic returns, $J^{a}(\pi_{\theta})$ in orange and $J^{s}(\pi_{\theta})$ in blue.
-The intrinsic reward $\rho^{a}(s,a)=-\log\pi_{\theta}(a|s)$, from equation (4), and the intrinsic reward $\rho^{s}(s,a)=-\log d^{\pi_{\theta},\gamma}(s)$, from equation (3), are used respectively.
+The intrinsic reward $\rho^{a}(s,a)=-\log\pi_{\theta}(a|s)$, from equation (3), and the intrinsic reward $\rho^{s}(s,a)=-\log d^{\pi_{\theta},\gamma}(s)$, from equation (4) are used, respectively.
 In Figure 3(b), we illustrate the return of the policy without exploration $J(\pi_{\theta})$, along with two learning objective functions, $L^{a}(\theta)$ and $L^{s}(\theta)$, using as exploration strategies the intrinsic returns $J^{a}(\pi_{\theta})$ and $J^{s}(\pi_{\theta})$.
 We observe that the return is a pseudoconcave function with respect to $\theta$ and the optimal parameter is $\theta^{*}=1$.
-In addition, the two learning objectives respect the $\epsilon$-coherence criterion for $\epsilon=0$, implying that $\theta^{*}=\theta^{\dagger}$, and respect the pseudoconcavity criterion.
+In addition, the two learning objectives satisfy the $\epsilon$-coherence criterion for $\epsilon=0$, implying that $\theta^{*}=\theta^{\dagger}$, and also satisfy the pseudoconcavity criterion.
 It is important to note that with regard to the discussion from Section 3, there is no interest in optimizing the learning objectives rather than directly optimizing the return, as the latter is already pseudoconcave.
-In the following we illustrate how choosing a correct exploration strategy still deeply influences the policy-gradient algorithms when it comes to building gradient estimates.
+In the following, we illustrate how choosing a correct exploration strategy still deeply influences the policy-gradient algorithms when it comes to building gradient estimates.
 (a) Return functions.
@@
 These intervals represent the smallest balls containing the parameters maximizing the intrinsic return and the learning objective, for both exploration strategies.
-Let us compute the estimate $\hat{g}$ and $\hat{d}$ relying on REINFORCE (Williams, 1992) by sampling $8$ histories of length $T=100$.
-In this particular environment, $\mathbb{P}(D>0)$ equals $\mathbb{P}(G>0)$, and equal the probability that the derivative is positive.
+Let us compute the estimates $\hat{g}$ and $\hat{d}$ relying on REINFORCE (Williams,, 1992) by sampling $8$ histories of length $T=100$.
+In this particular environment, $\mathbb{P}(D>0)$ equals $\mathbb{P}(G>0)$, and equals the probability that the derivative is positive.
 We represent in Figure 3(c) this probability for the return and for both learning objectives.
@@
 Let the minima of the orange and blue curves over these intervals be denoted by $\delta^{a}$ and $\delta^{s}$.
-By definition of the attraction criterion, it is thus respected for any values of $\delta$ at most equal to $\delta^{a}$ and $\delta^{s}$, for $L^{a}(\theta)$ and $L^{s}(\theta)$, respectively.
-All these observations can eventually be explained as the computation of $\hat{g}$ is always zero when the target is not sampled in the histories, which is highly likely for policies with small values of $\theta$.
+By definition of the attraction criterion, it is thus satisfied for any values of $\delta$ at most equal to $\delta^{a}$ and $\delta^{s}$, for $L^{a}(\theta)$ and $L^{s}(\theta)$, respectively.
+All these observations can eventually be explained because the computation of $\hat{g}$ is always zero when the target is not sampled in the histories, which is highly likely for policies with small values of $\theta$.
 Policy-gradient algorithms relying on intrinsic exploration would compute optimal policies efficiently where naive optimization without exploration would fail or be sample inefficient.
-We have empirically shown that a well-chosen exploration strategy in policy gradients may not only remove local extrema from the objective function, but may also increase the probability that stochastic ascent steps improve the objective function.
+We have empirically shown that a well-chosen exploration strategy in policy gradients may not only remove local extrema from the objective function, but may also increase the probability that stochastic ascent steps improve the objective function and eventually the return.
 Under the previous assumptions, this probability measures the efficiency of algorithms.
 Furthermore, among different learning objectives respecting the coherence and pseudoconcavity criteria, it is best to choose one that has high values for $\delta$ in both the efficiency and attraction criteria.
-In Appendix A we use these criteria to study other reward-shaping strategies, and in Appendix B we extend the study to more complex environments from the MiniGrid library (Chevalier-Boisvert et al., 2023) where the policy is a deep neural network.
+In Appendix A, we extend the study to more complex environments and policies.
 The problem discussed in this section strongly relates to overfitting or generalization in reinforcement learning.
@@
 Formally, a policy $\pi$ is locally optimal over a space with probability $\Delta$ if, and only if,
-[…]
-[…]
-In the typical case of environments with sparse rewards, many policies observe with high probability state and action pairs with zero rewards and are locally optimal for large probabilities $\Delta$.
-Typically, in the previous example, the joint set $\{1,\dots,S-2\}\times\{-1,1\}$ is a set of state and action pairs $\mathcal{E}$ that respects the definition equation (14) for large values $\Delta$ when $\theta$ is small.
+$\exists\mathcal{E}$ $\in\left\{\mathcal{X}\Big{|}\underset{\begin{subarray}{c}s\sim d^{\pi,\gamma}(\cdot)\\ a\sim\pi(\cdot|s)\end{subarray}}{\mathbb{E}}\big{[}\mathbbm{1}\left[(s,a)\in\mathcal{X}\right]\big{]}\geq\Delta\right\}:\pi\in\arg\max_{\pi^{\prime}}\underset{\begin{subarray}{c}s\sim d^{\pi^{\prime},\gamma}(\cdot)\\ a\sim\pi^{\prime}(\cdot|s)\end{subarray}}{\mathbb{E}}\big{[}\mathbbm{1}\left[(s,a)\in\mathcal{E}\right]\rho(s,a)\big{]}\;.$ (14)
+In the case of sparse rewards, many policies observe with high probability state and action pairs with zero rewards and are locally optimal for large probabilities $\Delta$.
+Typically, in the previous example, the joint set $\{1,\dots,S-2\}\times\{-1,1\}$ is a set of state and action pairs $\mathcal{E}$ that satisfies the definition in equation (14) for large values $\Delta$ when $\theta$ is small.
 As we have shown, exploration mitigates the convergence of policy-gradient algorithms towards these locally optimal policies.
 Note that assuming a non-zero reward is uniformly distributed over the state and action space, exploration policies with uniform probabilities over visited states and actions are the best choice for sampling non-zero rewards with high probability.
-It can thus also be considered as the best choice of exploration to reduce the probability that the stochastic gradient ascent steps do not increase the objective value.
-Such initial policy may be learned from the framework developed by Lee et al.
-(2019).
+It can thus also be considered the best choice of exploration to reduce the probability that the stochastic gradient ascent steps do not increase the objective value.
+Such initial policy may be learned as suggested by Lee et al., (2019).
 5 Conclusion
-In conclusion, this research takes a step towards dispelling misunderstandings about exploration through the study of its effects on the performance of policy-gradient algorithms.
-More particularly, we distinguished two effects exploration has on the optimization.
-First, it modifies the learning objective in order to remove local extrema.
+In conclusion, this research takes a step toward dispelling misunderstandings about exploration through the study of its effects on the performance of policy-gradient algorithms.
+More specifically, we distinguish two effects exploration has on the optimization.
+First, it modifies the learning objective to remove local extrema.
 Second, it modifies the gradient estimates and increases the likelihood that the update steps lead to improved returns.
-These two phenomena were studied through four criteria that we introduced and illustrated.
+These two phenomena are studied through four criteria that we introduce and illustrate.
 These ideas apply to other direct policy optimization algorithms.
 Indeed, the four criteria do not assume any structure on the learning objective and can thus be straightforwardly applied to any objective function optimized by a direct policy search algorithm.
-In particular, for off-policy policy gradient, we may simply consider that the off-policy objective is itself a surrogate or that the gradients of the return are biased estimates based on past histories.
+In particular, for off-policy policy-gradient methods, we may simply consider that the off-policy objective is itself a surrogate or that the gradients of the return are biased estimates based on past histories.
 Ideas introduced in this work also apply to other reinforcement learning techniques.
-Typically, for value-based RL with sparse-reward environments, convergence towards a value function outputting zero is expected with high probability.
-This is mostly due to the low probability of sampling non-zero rewards by Monte-Carlo.
-The discussions from Section 4 then apply, and a similar analysis can be performed.
-Our framework opens the door for further theoretical analysis, and the potential development of new criteria.
-We believe that deriving practical conditions on the exploration strategies, and the scheduling of the intrinsic return, for guaranteeing fast convergence should be the focus of attention.
-It could be achieved by bounding the policy improvement on expectation, which is nevertheless usually a hard task without strong assumptions.
-We furthermore believe that we provide a new lens on exploration necessary for interpreting and developing exploration strategies, in the sense of optimizing surrogate learning objective functions.
-Acknowledgments
-The authors thank Arnaud Delaunoy, Pascal Leroy, and Mathias Berger for valuable comments on this manuscript.
-Adrien Bolland and Gaspard Lambrechts gratefully acknowledge the financial support of the F.R.S.-FNRS.
-Appendix A Reward Shaping and Exploration Strategies
-As discussed in the manuscript, exploration strategies are reward-shaping strategies where the intrinsic reward bonuses are, among others, dependent on the policy parameters.
-This dependency makes the shaping strategies adaptive but makes the computation of gradients and the study of the learning objectives more complex.
-In this section, we study handcrafted reward-shaping strategies that have pseudoconcave and dense reward functions in the hill and maze environments.
-We then illustrate that the same criteria can be used to study these expert-knowledge based shaped rewards.
-For the hill environment from Section 3, we illustrate in Figure 4(a) an intrinsic reward bonus making the sum of rewards in equation (2) concave.
-The corresponding learning objective has a unique maximum, which is part of the set $\Omega=\left\{\theta^{\prime}|\max_{\theta}J(\pi_{\theta})-J(\pi_{\theta^{\prime}})\leq\epsilon\right\}$ with $\epsilon=1$ and $\theta=(K,\sigma)$.
-It can be seen in Figure 4(b) where the global maximum in black is within the set $\Omega$ in green.
-Both, the $\epsilon$-coherence and the pseudoconcavity criteria are thus respected for $\epsilon=1$.
-Here, the intrinsic reward function is a simple function independent of the policy $\pi_{\theta}$.
-Finding such an intrinsic reward may be complex for other environments but the example underlines that exploration and reward shaping are mostly equivalent and that designing reward functions that are concave may help converging towards optimal policies.
-(a) Alternative reward function.
-(b) Learning objective.
-Figure 4: In Figure 4(a), an alternative intrinsic reward function ensuring that the sum of rewards is a pseudoconcave function. In Figure 4(b), the contour function of the learning objective.
-For the maze environment, the return $J(\pi_{\theta})$ is represented in black in Figure 5(a) together with the intrinsic return $J^{d}(\pi_{\theta})$ in green.
-The latter is the return of the dense handcrafted reward function $\rho^{d}(s,a)=(a-1)/2$ penalizing actions moving away from the target.
-In Figure 5(b), the corresponding learning objective function is shown.
-In the same experimental setting as in Section 4, we observe that the objective function is $\delta$-efficient for higher values of $\delta$ compared to the already-discussed learning objectives.
-Furthermore, the attraction criterion is respected for any value of $\delta$ as the unique global maxima of the learning objective, intrinsic return, and return are all equals.
-(a) Return.
-(b) Learning objectives.
-(c) Improvement probabilities.
-Figure 5: In Figure 5(a) the return of the maze environment is represented together with the intrinsic return of a dense handcrafted reward function.
-Figure 5(b) represents the corresponding learning objective and Figure 5(c) the probability that the REINFORCE estimates are positive.
-Appendix B Minigrid Experiments
+Typically, for value-based RL with sparse-reward environments, convergence to a value function that outputs zero is expected with high probability.
+This is mostly due to the low probability of sampling non-zero rewards via Monte Carlo.
+Similar discussion and analysis as in Section 4 then apply.
+Our framework opens the door to further theoretical analysis, and the potential development of new criteria.
+We believe that deriving practical conditions on the exploration strategies and the scheduling of the intrinsic return to guarantee fast convergence should be the focus of attention.
+It could be achieved by bounding the policy improvement in expectation, which is nevertheless usually a hard task without strong assumptions.
+We further believe that we provide a new lens on exploration that is necessary for interpreting and developing exploration strategies, in the sense of optimizing surrogate learning objective functions.
+Appendix A Extended Experiments
 In this section, we introduce complex environments and parameterize policies with neural networks.
-In this context, it is impractical to naively compute and represent the objective functions and probability distributions for the different criteria.
-Therefore, we only evaluate the criteria along parameter trajectories, and extend the previous experimental setting.
-We consider seven environments from the MiniGrid suite of environments (Chevalier-Boisvert et al., 2023), among others, designed for evaluating exploration strategies.
-In these environments, an agent moves in a maze and aims to reach a target position.
-To do so, the agent may choose actions that consist of turning left, turning right, moving forward, or staying idle.
-We consider two reward settings: the dense setting and the sparse setting.
-In the first, rewards of $-1$ are received for every non-idle move, and a reward of $1000$ is received upon reaching the target position.
+We then learn the policy parameters using policy gradients and intrinsic exploration bonuses.
+We discuss convergence results based on the four criteria introduced in the paper.
+In this context, it is impractical to naively evaluate the different criteria for each policy parameter, and we are restricted to a local evaluation of the criteria on the basis of statistics computed during optimization.
+Let us adapt environments from the MiniGrid suite of environments (Chevalier-Boisvert et al.,, 2023).
+Each environment corresponds to a maze in which an agent moves, aiming to reach a target position.
+The agent may choose as actions to turn left, turn right, move forward, or stay idle.
+We consider two reward settings: dense and sparse.
+In the first, rewards of 1-1 are received for every non-idle move, and a reward of $1000$ is received upon reaching the target position.
 In the second setting, zero rewards are received everywhere, except upon reaching the target position, where a bonus of $1000$ is provided.
-In the dense setting, due to the action penalization incurred when moving, a policy outputting the idle action with probability one is locally optimal and has a return equal to zero.
-This is not (necessarily) the case in the second setting.
-We consider a discount factor of $\gamma=0.98$ and optimize a fully connected neural network taking as input the position pair and the orientation of the agent, and outputting a categorical distribution over actions.
+We consider $\gamma=0.98$.
+In the following, the policy is parameterized with a fully connected neural network taking as input the position pair and the orientation of the agent, and outputting a categorical distribution over actions.
 The network is composed of three hidden layers of $64$ neurons with ReLU activation functions.
-In the dense reward setting, we optimize policies by maximizing three learning objective functions: $J(\pi_{\theta})$, $L^{a}(\theta)$, and $L^{s}(\theta)$, respectively with $\lambda_{a}=0.5$ and $\lambda_{s}=0.25$.
-For the last objective, the state-visitation density estimator is a ten-component Gaussian mixture model maximizing the likelihood of the sampled batch.
-The optimization is performed using the Adam update rule (Kingma & Ba, 2014), with REINFORCE ascent directions computed over 32 histories of constant length $T=100$, and with learning rate (step size) equal to $0.0005$.
-The length $T$ of the histories is chosen such that the realization value $T$ from a geometric distribution with success probability parameter $1-\gamma$ has at least a cumulative probability of $0.85$.
-In this setting, we illustrate the quasiconcavity criterion and the $\epsilon$-coherence criterion.
+The policy is optimized in the different environments and reward settings.
+To that end, we consider three learning-objective functions: $J(\pi_{\theta})$, $L^{a}(\theta)$, and $L^{s}(\theta)$, respectively with $\lambda_{a}=0.5$ and $\lambda_{s}=0.25$.
+For the last objective, the state-visitation density estimator is a ten-component Gaussian mixture model that maximizes the likelihood of the sampled batch.
+The optimization is performed using the Adam update rule (Kingma and Ba,, 2014), with REINFORCE ascent directions computed over $32$ histories of constant length $T=100$, and with a learning rate (step size) equal to $0.0005$.
+The length $T$ of the histories is chosen such that the realized value of $T$ from a geometric distribution with success probability parameter $1-\gamma$ has at least a cumulative probability of $0.85$.
+We first consider the environments in the dense-reward setting and discuss convergence in light of the pseudoconcavity and $\epsilon$-coherence criteria.
 In Figure 4, we provide the evolution of the return of the policies when optimizing the three objectives $J(\pi_{\theta})$, $L^{a}(\theta)$, and $L^{s}(\theta)$ for the different environments.
-For the MiniGrid-Empty-8x8-v0 and the MiniGrid-FourRooms-v0 environments, optimizing the return results in high-performance policies that do not stay idle.
-The other objectives also manage to find high-performing policies, but with a lower return.
-This phenomenon (assuming that the global optimum of each objective is found) illustrates the $\epsilon$-coherence criterion, where this $\epsilon$ value is the bound on the best policy that can be found when optimizing the learning objective.
-For the other environments, the policies resulting from the optimization of the return fall into local optima, namely ones where the policy chooses the idling action with probability one.
-When optimizing the learning objectives with exploration bonuses, the resulting policies no longer fall into the previous local optima.
-This result suggests that, along these parameter trajectories, the return $J(\pi_{\theta})$ has a local optimum (or saddle point), in opposition to the learning objective functions $L^{a}(\theta)$ and $L^{s}(\theta)$.
+On the one hand, in the MiniGrid-Empty-8x8-v0 and MiniGrid-FourRooms-v0 environments, optimizing the return provides high-performing policies, and the policies resulting from the optimization of the other learning objectives are suboptimal.
+The difference in performance is justified by the $\epsilon$-coherence criterion, where $\epsilon$ is the bound on the best policy that can be found when optimizing the learning objective (assuming that the global optimum of each objective is found).
+On the other hand, in the other environments, optimizing the return often leads to convergence to locally optimal parameters (or saddle points).
+These parameters correspond to idling policies and are locally optimal due to the action penalization incurred when moving.
+The policies obtained when optimizing $L^{a}(\theta)$ and $L^{s}(\theta)$ no longer fall into the previous local optima.
 The latter illustrates the validity of the pseudoconcavity criterion in that region of the parameter space.
-For the learning objective $L^{s}(\theta)$, the $\epsilon$-coherence criterion is respected for a small value of $\epsilon$, and the resulting policy manages to reach the target position.
-For the learning objective $L^{a}(\theta)$, the resulting policy does not reach the target position.
-We nevertheless hypothesize that it is not a real local optimum of the learning objective but a local optimum in the sense of equation (14).
-Figure 6: Evolution of the return of policies during optimization in the dense minigrid environments.
+Note that when optimizing $L^{a}(\theta)$, the policy tends to converge towards another local optimum where actions are uniformly distributed.
+We nevertheless hypothesize it is local in the sense of equation (14).
+Figure 4: Return of policies during optimization in the dense minigrid environments.
 In blue, the return $J(\pi_{\theta})$ is optimized; in orange, the learning objective $L^{a}(\theta)$ is optimized; and in green, the learning objective $L^{s}(\theta)$ is optimized by performing Adam steps in REINFORCE directions.
 Note that the median, worst, and best cases over five runs are represented for the different curves.
-For the environments MiniGrid-Empty-8x8-v0 and MiniGrid-FourRooms-v0, optimizing each objective results in a policy that does not stay idle.
-These policies are initialized outside of the basin of attraction of the local optimum of the return.
-The coherence criteria can be observed as optimizing the learning objective with intrinsic exploration bonuses results in suboptimal policies.
-For the other environments, optimizing the return directly leads to policies that always choose to stay idle, and are thus locally optimal.
-Optimizing the learning objective with exploration allows us to escape from these local optima, illustrating the quasiconcavity criterion.
-It can be noted that when optimizing the objective $L^{a}(\theta)$, the $\epsilon$-coherence criterion is respected for a large value of $\epsilon$, making the resulting policy worse than when optimizing the return directly.
-On the contrary, the objective $L^{s}(\theta)$ appears to have $\epsilon$-coherence for a reasonable value of $\epsilon$.
-In the previous experiments with the dense setting, the local optima exist due to the negative rewards associated to idle-actions.
-If we consider the sparse setting, we could then assume that directly optimizing the return is sufficient to find high-performing policies.
-It is not always the case and it can be justified using the efficiency and attraction criteria.
-We use the same parameters as in the previous set of experiments and we provide the evolution of the return of the policies when optimizing the three objectives $J(\pi_{\theta})$, $L^{a}(\theta)$, and $L^{s}(\theta)$, for the different environments in Figure 7.
-On the one hand, for most environments, whatever the learning objective, the resulting policy has a high return.
-Note that the $\epsilon$-coherence can again be illustrated where the policies resulting from the optimization of the return perform better than the others.
-On the other hand, for the MiniGrid-Empty-16x16-v0 and the MiniGrid-SimpleCrossingS11N5-v0 environments, a high performing policy can only be found when optimizing the learning objective $L^{s}(\theta)$.
-We illustrate that these results can be justified by the efficiency and attraction criteria in Figure 8.
-For each parameters obtained during the stochastic ascent steps on the return, we first estimate the probability of improving both objective functions by stochastic gradient ascent.
-These probabilities are used to compare learning objectives in terms of the efficiency criterion; the larger, the better.
-Then, in order to illustrate the attraction criterion, we estimate the probability of improving the return and the learning objective, both by gradient ascent steps on the learning objective, for each parameters obtained during the stochastic ascent steps on the objective $L^{a}(\theta)$ and $L^{s}(\theta)$.
-As far as $L^{a}(\theta)$ is concerned, all probabilities remain small as result that the optimization procedure converges fast towards a stationary point where the target goal is observed with negligible probability.
-The efficiency and attraction criteria are respected for negligible probabilities $\delta$, which is also a justification for the failure of converging towards good policies in the dense setting.
-For $L^{s}(\theta)$, on the contrary, and in both environments, the probability of improving the learning objective remains large for each parameter encountered when optimizing the return.
-The efficiency of the learning objective is much higher than that of the return in that part of the parameter space.
-Furthermore, the probability of improving the return when optimizing the learning objective, is small at the beginning and increases after some iterations.
-This indicates that once the policy has a sufficiently large intrinsic return, the attraction criterion is respected for a high value $\delta$.
-Figure 7: Evolution of the return of policies during optimization in the sparse minigrid environments.
+In the previous experiments, the local optima arise due to the negative rewards associated with idle actions.
+If we consider the sparse-reward setting, we could then assume that directly optimizing the return is sufficient to find high-performing policies.
+We use the same parameters as previously and provide the evolution of the return of the policies when optimizing the three objectives $J(\pi_{\theta})$, $L^{a}(\theta)$, and $L^{s}(\theta)$ for the different environments in Figure 5.
+On the one hand, for most environments, regardless of the learning objective, the resulting policy has a high return.
+Note that the $\epsilon$-coherence can again be illustrated, as the policies resulting from the optimization of the return perform better than the others.
+On the other hand, for the MiniGrid-Empty-16x16-v0 and MiniGrid-SimpleCrossingS11N5-v0 environments, a high-performing policy can only be found when optimizing the learning objective $L^{s}(\theta)$.
+We illustrate in the following that these results can be justified by the efficiency and attraction criteria.
+Figure 5: Return of policies during optimization in the sparse minigrid environments.
 In blue, the return $J(\pi_{\theta})$ is optimized; in orange, the learning objective $L^{a}(\theta)$ is optimized; and in green, the learning objective $L^{s}(\theta)$ is optimized performing Adam steps in REINFORCE directions.
 Note that the median, worst, and best cases over five runs are represented for the different curves.
-In most environments, a high-performing policy can be found by optimizing the return $J(\pi_{\theta})$.
-This results from its quasiconcavity.
-The $\epsilon$-coherence criterion can also be observed.
-However, for the MiniGrid-Empty-16x16-v0 and the MiniGrid-SimpleCrossingS11N5-v0 environments, a high-performing policy can only be found when optimizing the learning objective $L^{s}(\theta)$.
-Figure 8: For the MiniGrid-Empty-16x16-v0 and the MiniGrid-SimpleCrossingS11N5-v0 environments, and for both learning objective functions $L^{a}(\theta)$ and $L^{s}(\theta)$, we first represent the estimated probability of improving the return $J(\pi_{\theta})$ and the corresponding learning objective when following their REINFORCE gradient estimate.
-This value is estimated at each run of the optimization of the policy with learning objective $J(\pi_{\theta})$.
-Second, we represent the estimated probability of improving the corresponding learning objective and the return $J(\pi_{\theta})$ when following the REINFORCE gradient estimate of the learning objectives.
-These values are estimated at each run of the optimization of the policy with the learning objectives.
-The probabilities were estimated based on the frequencies of improving the objective functions by more than $0.2$ when following $5$ Adam ascent steps using REINFORCE update directions.
+In Figure 6(a), we show the probability of improving the return $J(\pi_{\theta})$ and the learning objective $L^{s}(\theta)$ for each parameter obtained during their respective optimization by stochastic gradient ascent steps (i.e., SGA following respectively $\hat{g}\approx\nabla_{\theta}J(\pi_{\theta})$ and $\hat{d}\approx\nabla_{\theta}L^{s}(\theta)$).
+The improvement probability is higher for the learning objective than for the return.
+In other words, for these parameters, the efficiency $\delta$ of the learning objective is much higher than that of the return, which justifies the failure of optimization without exploration in Figure 5.
+In Figure 6(b), we again show the probability of improving the return $J(\pi_{\theta})$ and the learning objective $L^{s}(\theta)$, but for each parameter obtained during the stochastic gradient ascent steps on $L^{s}(\theta)$ (i.e., SGA following $\hat{d}\approx\nabla_{\theta}L^{s}(\theta)$).
+The probability of improving the return when optimizing the learning objective is small at the beginning and increases after some iterations.
+This indicates that after some iterations, the attraction criterion holds for a large $\delta$.
+The return thus eventually starts increasing, and optimizing $L^{s}(\theta)$ provides a high-performing policy, as can be seen in Figure 5.
+(a) SGA following $\hat{g}$ and $\hat{d}$
+(b) SGA following $\hat{d}$
+Figure 6: In Figure 6(a), we show the estimated probability of improving the return $J(\pi_{\theta})$ and the learning objective $L^{s}(\theta)$ when following their REINFORCE gradient estimates.
+In Figure 6(b), we show the estimated probability of improving the return $J(\pi_{\theta})$ and the learning objective $L^{s}(\theta)$ when following the REINFORCE gradient estimate of the learning objective $L^{s}(\theta)$.
+Probabilities were estimated based on the frequencies of improving the objective functions by more than $0.2$ after following 55 Adam ascent steps.